### v1

# Control variates for variance-reduced ratio of means estimators

Control variates for variance-reduced ratio of means estimators
Louison Bocquet-Nouaille Corresponding author
Emails: louison.bocquet_nouaille@onera.fr, jerome.morio@onera.fr, benjamin.bobbia@isae-supaero.fr ONERA/DTIS, Université de Toulouse, F-31055 Toulouse Fédération ENAC ISAE-SUPAERO ONERA, Université de Toulouse, 31000 Toulouse
Jérôme Morio ONERA/DTIS, Université de Toulouse, F-31055 Toulouse Fédération ENAC ISAE-SUPAERO ONERA, Université de Toulouse, 31000 Toulouse Benjamin Bobbia Fédération ENAC ISAE-SUPAERO ONERA, Université de Toulouse, 31000 Toulouse
( Preprint - October 15, 2025)
The control variates method is a classical variance reduction technique for Monte Carlo estimators that exploits correlated auxiliary variables without introducing bias. In many applications, the quantity of interest can be expressed as a ratio of expectations. We propose a variance-reduced estimator for such ratios, which applies control variates to both the numerator and the denominator. The control variate coefficients are optimized jointly to minimize the variance of the resulting estimator. This approach theoretically guarantees variance reduction and naturally extends to approximate control variates. Simulation studies show significant variance reduction, particularly when correlations between variables and control variates are strong. The practical value of the method is illustrated with a multi-fidelity aircraft design use case.
1 Introduction
The variance of the Monte Carlo empirical mean estimator is inversely proportional to the number of available samples, which is often limited in real-world applications. As a result, every observation has a large influence on the estimate, leading to high variance and reduced reliability.
An overview of common variance reduction techniques can be found in [Asmussen and Glynn, 2007]. In the setting considered here, where only a limited number of samples is available and further sampling is not possible, most methods are not directly applicable. The control variates method [Asmussen and Glynn, 2007, Sec. 2] is well-suited, as it exploits i.i.d. samples from an auxiliary variable correlated with the target variable, to reduce the variance of the Monte Carlo estimator without introducing bias. It can be seen as a form of transfer learning [Zhu et al., 2025], where information from a source variable is used to improve estimation on the target variable.
One of the first references to the control variates method appears in [Hammersley and Handscomb, 1965], Section 2.6, where it is introduced under the name of “regression estimator.” The method can be interpreted as a linear regression, where the coefficient of the control variate is chosen as the slope that best explains the variability of the output through its correlation with the auxiliary variable. Later works developed this regression view further: [Lewis et al., 1987] with linear and nonlinear adjustments, [Davidson and MacKinnon, 1992] with applications to quantile estimation, and [Salaün et al., 2022] with regression-based Monte Carlo integration. Control variates have been adapted in many directions, from theoretical analyses to practical implementations; in the following, we mention a selection of approaches. [Hesterberg and Nelson, 1998] discusses the choice of control variates when estimating probabilities or quantiles. [Glynn and Szechtman, 2002] provides a comprehensive Hilbert-space perspective on control variates, establishes asymptotic optimality results, and explores their connections with other variance reduction techniques. [Hickernell et al., 2005] extends the control variates method to quasi-Monte Carlo estimators, highlighting that the optimal control variate differs from the classical Monte Carlo setting. [Müller et al., 2020] introduces Neural Control Variates, using a set of neural networks to apply control variates to parametric Monte Carlo integration. [Fox and Ökten, 2021] presents a polynomial chaos expansion method for control variates. [Maskell et al., 2022] extends the method to the estimation of constrained variables.
The approximate control variates method extends the standard control variates approach to cases where the mean of the auxiliary variable is unknown, estimating it from additional samples. It is also referred to as Multi-Fidelity Monte Carlo (MFMC) [Cannamela et al., 2008] [Gorodetsky et al., 2020] [Kim et al., 2024]: the output of a computationally cheaper lower-fidelity model is used as a control variate to improve the estimation on the output of a more accurate high-fidelity model. Since both models approximate the same underlying quantity, their outputs are often highly correlated, which enables efficient variance reduction through control variates.
In this work, we focus on ratio of means estimation, a quantity of interest in many applications. The standard approach consists of estimating the numerator and denominator separately by Monte Carlo, and then taking their ratio. This estimator has been analysed in [Cochran, 1977]. The control variates approach can be applied to both components to reduce variance. To the best of our knowledge, [Gordon et al., 1982] is the only work specifically extending the method to ratio of means estimation, by setting the numerator coefficient to its classical value, the one minimizing variance for a single mean, and optimizing the denominator coefficient to minimize the ratio variance.
The classical coefficients for mean estimation, or those of [Gordon et al., 1982], can in fact increase the variance of the ratio of means estimator with certain correlation structures, as they are not both derived to minimize the variance of the ratio. This paper introduces new optimal coefficients that theoretically guarantee variance reduction as they are jointly optimized to minimize the variance of the estimator.
The paper is organized as follows. In Section 2, the control variates method is presented for mean estimation, both in its exact and approximate forms, and for ratio of means estimation. Section 3 introduces the new optimal coefficients for variance-reduced ratio of means estimators using control variates. In Section 4, simulation results are presented, illustrating extreme correlation structures and their impact on variance reduction. Finally, in Section 5, the proposed estimators are applied to a multi-fidelity aircraft design dataset.
2 Variance reduction with control variates
2.1 Control variates applied to a mean estimator
The control variates method aims to reduce the variance of the Monte Carlo estimator defined as follows.
Definition 1 (Monte Carlo estimator).
Let $(A_{i})_{i=1\dots n}$ be i.i.d. real samples of a random variable $A\in L^{1}$. The Monte Carlo estimator $\overline{A_{n}}$ of $\mathbb{E}[A]$ is defined as
$\overline{A_{n}}=\frac{1}{n}\sum_{i=1}^{n}A_{i}$ (1)
The variance of the Monte Carlo estimator is proportional to the number of samples $n$:
$\text{Var}(\overline{A_{n}})=\frac{1}{n}\text{Var}(A).$ (2)
To reduce the variance without increasing $n$, the idea is to consider a control variate $B$, whose expectation $\mathbb{E}[B]$ is known and correlated with the target random variable $A$. This approach is a form of statistical transfer learning: it leverages knowledge from a related variable $B$ to compensate for the lack of information about the target variable $A$. The control variates estimator of $\mathbb{E}[A]$ is defined as follows.
Definition 2 (Control variates estimator [Asmussen and Glynn, 2007]).
Let $(A_{i},B_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$ and $B$. Assume $\mathbb{E}[B]$ is known. The control variates estimator of $\mathbb{E}[A]$ is defined as
$\widehat{A}_{CV}=\overline{A_{n}}+\alpha_{c}(\mathbb{E}[B]-\overline{B_{n}})$ (3)
where the coefficient $\alpha_{c}$ is set to minimize the variance of $\widehat{A}_{CV}$:
$\alpha_{c}:=\underset{\alpha\in\mathbb{R}}{\text{argmin}}\;\text{Var}\left(\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})\right)=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ (4)
The coefficient $\alpha_{c}$ needs to be estimated, which can introduce a small bias if the estimation is made from the same dataset used to compute $\widehat{A}_{CV}$. This plug-in bias [Kennedy, 2024] is generally negligible in practice, as discussed in [Owen, 2013].
With the optimal coefficient $\alpha_{c}$, the variance of the control variates estimator is
$\text{Var}(\widehat{A}_{CV})=\left(1-\text{Corr}(A,B)^{2}\right)\text{Var}(\overline{A_{n}})$ (5)
where $\text{Corr}(A,B)$ is the linear correlation coefficient between $A$ and $B$. This expression shows that the method never increases variance. The extent of variance reduction depends on the strength of the correlation between $A$ and $B$: the higher the absolute value of the correlation, the greater the reduction. When the correlation is zero, the control variates estimator coincides with the Monte Carlo one.
In practice, the assumption that the control variate mean $\mathbb{E}[B]$ is known is often unrealistic. The approximate control variates (ACV) method overcomes this by estimating the mean from additional data.
Definition 3 (Approximate control variates [Gorodetsky et al., 2020]).
Let $(A_{i},B_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$ and $B$, and let $(B_{i})_{i=n+1\dots n+m}$ be additional i.i.d. samples of $B$. The approximate control variates estimator of $\mathbb{E}[A]$ is defined as
$\widehat{A}_{ACV}=\overline{A_{n}}+\alpha_{c}(\overline{B_{n+m}}-\overline{B_{n}})$ (6)
The variance of the approximate control variates estimator is
$\text{Var}(\widehat{A}_{ACV})=\left(1-\frac{m}{n+m}\text{Corr}(A,B)^{2}\right)\text{Var}(\overline{A_{n}}).$ (7)
[…]
The control variate $B$ is typically chosen to mirror the variable it controls [Hesterberg and Nelson, 1998].
While the method naturally extends to multiple control variates using linear regression to estimate the various coefficients [Leluc et al., 2021], only a single control variate is considered here for clarity.
2.2 Control variates applied to ratio of means estimators
This section presents a variance-reduced estimator based on control variates, for the ratio of means $R$ defined as
$R=\frac{\mathbb{E}[A]}{\mathbb{E}[C]}$ (8)
where $A$ and $C$ are arbitrary random variables. Assume $\mathbb{E}[C]\neq 0$. The control variates method can be applied to reduce the variance of the Monte Carlo ratio of means estimator
$\widehat{R}_{\frac{MC}{MC}}=\frac{\overline{A_{n}}}{\overline{C_{n}}}$ (9)
referred to as the MC/MC estimator, where $\overline{A_{n}}$ and $\overline{C_{n}}$ are the Monte Carlo estimators of $\mathbb{E}[A]$ and $\mathbb{E}[C]$, as in (1).
To reduce the variance of the MC/MC estimator, the control variates method can be applied to both the numerator and the denominator. We refer to the resulting estimator as the CV/CV estimator.
Definition 4 (CV/CV estimator).
Let $(A_{i},B_{i},C_{i},D_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$, $B$, $C$, and $D$, where $B$ and $D$ are control variates. The CV/CV estimator of $R$ is defined as
$\widehat{R}_{\frac{CV}{CV}}=\frac{\widehat{A}_{CV}}{\widehat{C}_{CV}}=\frac{\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})}{\overline{C_{n}}+\beta(\mathbb{E}[D]-\overline{D_{n}})}$ (10)
Proposition 1 (Coefficients for the CV/CV estimator).
Different choices of coefficients exist:
  • the classical coefficients for mean estimators, introduced in Definition 2:
    $\alpha_{c}:=\underset{\alpha\in\mathbb{R}}{\text{argmin}}\;\text{Var}\left(\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})\right)=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ (11)
    $\beta_{c}:=\underset{\beta\in\mathbb{R}}{\text{argmin}}\;\text{Var}\left(\overline{C_{n}}+\beta(\mathbb{E}[D]-\overline{D_{n}})\right)=\frac{\text{Cov}(C,D)}{\text{Var}(D)}$ (12)
  • the coefficients introduced in [Gordon et al., 1982]:
    $\alpha_{g}$ $:=\underset{\alpha\in\mathbb{R}}{\text{argmin}}\;\text{Var}\left(\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})\right)$ $=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ (13)
    $\beta_{g}$ $:=\underset{\beta\in\mathbb{R}}{\text{argmin}}\;\text{Var}\left(\frac{\overline{A_{n}}+\alpha_{g}(\mathbb{E}[B]-\overline{B_{n}})}{\overline{C_{n}}+\beta(\mathbb{E}[D]-\overline{D_{n}})}\right)$ $=\frac{R\text{Cov}(C,D)-\text{Cov}(A,D)+\alpha_{g}\text{Cov}(B,D)}{R\text{Var}(D)}$ (14)
However, these choices are not always optimal: variance reduction is not guaranteed in all cases, and the variance can even be increased as illustrated analytically in A.3 and with simulations in Section 4. The variance difference between the CV/CV and MC/MC estimators can be explicitly expressed, but analysing the conditions on the covariances and variances of $A$, $B$, $C$, and $D$ that ensure variance reduction remains a delicate task.
3 Optimized coefficients for control variates ratio of means estimators
3.1 Control variates applied to both numerator and denominator
In this section, we propose a new set of optimal coefficients that guarantee variance reduction with the CV/CV estimator.
Proposition 2 (Optimal coefficients for the CV/CV estimator).
Assume $|\text{Corr}(B,D)|<1$. The optimal coefficients for the CV/CV estimator, guaranteeing variance reduction, are defined as
$(\alpha_{o},\beta_{o}):=\underset{(\alpha,\beta)\in\mathbb{R}^{2}}{\text{argmin}}\;\text{Var}\left(\frac{\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})}{\overline{C_{n}}+\beta(\mathbb{E}[D]-\overline{D_{n}})}\right)$
Their analytical expression is given by
$\alpha_{o}=\frac{\text{Var}(D)\text{Cov}(A,B)-R\text{Var}(D)\text{Cov}(B,C)+R\text{Cov}(B,D)\text{Cov}(C,D)-\text{Cov}(B,D)\text{Cov}(A,D)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}$ (15)
$\beta_{o}=\frac{\text{Cov}(B,D)\text{Cov}(A,B)-R\text{Cov}(B,D)\text{Cov}(B,C)+R\text{Var}(B)\text{Cov}(C,D)-\text{Var}(B)\text{Cov}(A,D)}{R\left(\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}\right)}$ (16)
A.3 presents the full derivation of the optimal coefficients $\alpha_{o}$ and $\beta_{o}$. It is based on an approximate expression of the CV/CV estimator’s variance, which is shown to be a convex function of $(\alpha,\beta)$. The analytical formulas of the optimal coefficients are then obtained by setting the gradient of this variance to zero and solving for $(\alpha,\beta)$.
Proposition 3.
With the optimal coefficients given in (15) and (16), the variance reduction obtained with the CV/CV estimator is:
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$ $\approx-\frac{1}{n\mathbb{E}[C]^{2}}\frac{Var\bigg(\big(R\text{Cov}(B,C)-\text{Cov}(A,B)\big)D-\big(R\text{Cov}(C,D)-\text{Cov}(A,D)\big)B\bigg)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}\leq 0$ (17)
The variance reduction is guaranteed:
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)\leq\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$ (18)
If $|\text{Corr}(B,D)|<1$ and with the optimal coefficients presented here, the CV/CV estimator has a theoretical guarantee of variance reduction, as the control variates estimator for a simple mean.
From the variance difference formula given in (17), it seems the variance reduction will be greater when the variables $A$ and $C$ are correlated to the control variates $B$ and $D$. It is delicate to determine which covariances and variances should be maximized or minimized to boost variance reduction, as it depends on their signs, and many different correlation structures lead to variance reduction.
The mild condition $|\text{Corr}(B,D)|<1$, which ensures that the control variates are not perfectly linearly correlated, can be easily satisfied in practice. While it is possible to use two linearly correlated (or even identical) control variates, in such cases it is no longer possible to derive coefficients that guarantee variance reduction unconditionally.
Proposition 4.
Assume $\forall(a,b)\in\mathbb{R}^{*}\times\mathbb{R},B=aD+b$. The optimal coefficients for the CV/CV estimator with linearly correlated control variates $B$ and $D$ are
$\begin{cases}\alpha_{o}=\frac{1}{a}\left(\frac{\text{Cov}(A,D)-R\text{Cov}(C,D)}{\text{Var}(D)}+\beta_{o}R\right)\\ \beta_{o}\in\mathbb{R}\end{cases}$ (19)
Hence, there is variance reduction if
$\text{Cov}(A-RC,D)\in\left(-\sqrt{\frac{\text{Var}(D)}{2}},0\right)\cup\left(\sqrt{\frac{\text{Var}(D)}{2}},+\infty\right)$ (20)
The previous condition can be checked to determine whether the CV/CV estimator achieves variance reduction. Details of the derivation are given in A.3.3.
3.2 Control variates applied to the numerator only
An alternative strategy to derive optimal coefficients for the CV/CV estimator is to focus on the numerator only. When the denominator variance is already small, reducing it may be unnecessary. In such cases, the denominator coefficient $\beta$ can be set to zero, and only the numerator coefficient $\alpha$ is optimized. The resulting estimator is called the CV/MC estimator, as the control variates method is applied only to the numerator, while the denominator is estimated by Monte Carlo.
Definition 5 (CV/MC estimator).
Let $(A_{i},B_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$, $B$, and $C$, and let $(B_{i})_{i=n+1\dots n+m}$ be additional i.i.d. samples of $B$. The CV/MC estimator of $R$ is defined as:
$\widehat{R}_{\frac{CV}{MC}}=\frac{\widehat{A}_{CV}}{\overline{C_{n}}}=\frac{\overline{A_{n}}+\alpha_{o}^{\prime}(\mathbb{E}[B]-\overline{B_{n}})}{\overline{C_{n}}}$ (21)
where the coefficient $\alpha_{o}^{\prime}$ is set to minimize the variance of the ratio as
$\alpha_{o}^{\prime}:=\underset{\alpha\in\mathbb{R}}{\text{argmin}}\;\text{Var}\left(\frac{\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})}{\overline{C_{n}}}\right)=\frac{\text{Cov}(A,B)-R\text{Cov}(B,C)}{\text{Var}(B)}$ (22)
A.2 details the derivation of the optimal coefficient $\alpha_{o}^{\prime}$.
Proposition 5.
With the optimal coefficient $\alpha_{o}^{\prime}$ given in (22), the variance reduction obtained with the CV/MC estimator is:
$\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\approx-\frac{1}{n\mathbb{E}[C]^{2}}\frac{\left(\text{Cov}(A,B)-R\text{Cov}(B,C)\right)^{2}}{\text{Var}(B)}\leq 0$ (23)
The variance reduction is guaranteed:
$\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)\leq\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$ (24)
As the control variates mean estimator, the CV/MC ratio of means estimator presented here guarantees a variance no greater than that of the MC/MC estimator.
The CV/MC estimator requires only the numerator coefficient $\alpha_{o}^{\prime}$, which is simpler to estimate than the full pair $(\alpha_{o},\beta_{o})$ of the CV/CV estimator. Introducing the denominator coefficient $\beta$ adds an additional degree of freedom, allowing for further variance reduction at the cost of increased estimation complexity. Theoretically, the CV/CV estimator achieves a lower variance than the CV/MC estimator, as it minimizes the variance with respect to both $\alpha$ and $\beta$, whereas the CV/MC estimator restricts $\beta$ to zero. In practice, the CV/MC estimator may only perform comparably when the optimal $\beta_{o}$ is close to zero. Simulation results in Section 4.1 confirm that the CV/CV estimator consistently yields lower variance than the CV/MC estimator.
The MC/MC estimator $\widehat{R}_{\frac{MC}{MC}}$ is asymptotically unbiased as $n\to\infty$ [Cochran, 1977], a property that extends to the control variates estimators $\widehat{R}_{\frac{CV}{MC}}$ and $\widehat{R}_{\frac{CV}{CV}}$ introduced previously. For smaller sample sizes, bias may arise due to the nonlinearity of the ratio form: the expectation of a ratio is generally not equal to the ratio of expectations. This bias is particularly relevant when the denominator has high variance or is near zero. Nonetheless, as illustrated in Sections 4 and 5, the impact of this bias is negligible.
3.3 Approximate control variates for a ratio of means estimator
The CV/MC and CV/CV estimators with optimal coefficients previously introduced can be easily extended to the approximate control variates method, by replacing the control variates means $\mathbb{E}[B]$ and $\mathbb{E}[D]$ by Monte Carlo estimations on $n+m$ points:
$\widehat{R}_{\frac{ACV}{MC}}=\frac{\overline{A_{n}}+\alpha_{o}(\overline{B_{n+m}}-\overline{B_{n}})}{\overline{C_{n}}}$ (25)
$\widehat{R}_{\frac{ACV}{ACV}}=\frac{\overline{A_{n}}+\alpha_{o}(\overline{B_{n+m}}-\overline{B_{n}})}{\overline{C_{n}}+\beta_{o}(D_{n+m}-\overline{D_{n}})}$ (26)
As illustrated in A.4, the optimal coefficients do not change, and the theoretical variance reduction formulas are easily derived: they are the same as in the exact case with the additional scaling factor $\frac{m}{n+m}$:
$\text{Var}\left(\widehat{R}_{\frac{ACV}{MC}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)=\frac{m}{n+m}\left(\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\right)$ (27)
$\text{Var}\left(\widehat{R}_{\frac{ACV}{ACV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)=\frac{m}{n+m}\left(\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\right)$ (28)
4 Simulations on an academical use case
The reported simulation values are averaged over 10,000 independent repetitions, comparing different coefficients. For the CV/MC estimator, two coefficients are compared:
  • $\alpha_{c}$ the "classical" coefficient, given in (4), derived by optimizing the numerator only;
  • $\alpha_{o}^{\prime}$ the new optimal coefficient, given in (22), that minimizes the variance of the full ratio.
For the CV/CV estimator, three sets of coefficients are compared:
  • $(\alpha_{c},\beta_{c})$ the "classical" coefficients, given in (11) and (12), derived by optimizing the numerator and denominator independently;
  • $(\alpha_{g},\beta_{g})$ the coefficients proposed by [Gordon et al., 1982], given in (13) and (14), where $\alpha_{g}$ minimizes the numerator variance and $\beta_{g}$ minimizes the variance of the ratio;
  • $(\alpha_{o},\beta_{o})$ the new optimal coefficients, given in (15) and (16), that jointly minimize the variance of the full ratio.
Performance is measured with the relative variance reduction (RVR), that quantifies the variance reduction achieved by a new estimator $\widehat{R}_{\text{new}}$ compared to a baseline estimator $\widehat{R}_{\text{base}}$:
$\text{RVR}:=\frac{\text{Var}(\widehat{R}_{\text{base}})-\text{Var}(\widehat{R}_{\text{new}})}{\text{Var}(\widehat{R}_{\text{base}})}$ (29)
Here, the baseline $\widehat{R}_{\text{base}}$ denotes the MC/MC estimator. A positive RVR indicates that the new estimator $\widehat{R}_{\text{new}}$ achieves a lower variance than the baseline, while a negative RVR implies a variance increase.
The random variables $A,B,C,$ and $D$ are modelled as components of a four-dimensional Gaussian vector:
$(A,B,C,D)^{\top}\sim\mathcal{N}\left(\mu,\Sigma\right)\quad\text{with }\mu=\begin{pmatrix}50\\ 20\\ 10\\ 100\end{pmatrix}\quad\Sigma=\begin{pmatrix}1&\text{Cov}(A,B)&\text{Cov}(A,C)&\text{Cov}(A,D)\\ \text{Cov}(A,B)&1&\text{Cov}(B,C)&\text{Cov}(B,D)\\ \text{Cov}(A,C)&\text{Cov}(B,C)&1&\text{Cov}(C,D)\\ \text{Cov}(A,D)&\text{Cov}(B,D)&\text{Cov}(C,D)&1\end{pmatrix}$ (30)
With these parameters, the ratio $R=\mathbb{E}[A]/\mathbb{E}[C]$ to be estimated equals 5. All variances are normalized to one, so the covariance matrix coincides with the linear correlation matrix. For clarity, only the lower triangle of the covariance matrix is reported in each experiment. Unless otherwise specified, the sample size $n$ is 100.
Three particular configurations are considered in the following to illustrate possible behaviours of the estimators depending on the control variates coefficients. The first corresponds to the best-case scenario for the optimal coefficients proposed in this work, i.e. the covariance structure that leads to the largest variance reduction. The second corresponds to the worst-case scenario for the coefficients of [Gordon et al., 1982], i.e. the covariance structure that leads to the largest variance increase. The last one corresponds to the best-case scenario for the coefficients of [Gordon et al., 1982]. The covariance configurations are obtained by maximizing or minimizing the variance reduction using the formulas in A.2. The optimization is carried out using the differential evolution algorithm [Storn and Price, 1997], with the objective function set to infinity whenever the tested covariance matrix is not positive definite.
4.1 Strong variance reduction
The presented results are the best-case scenario for the proposed optimal coefficients, observed with the following covariance matrix
$\Sigma=\begin{pmatrix}1&&&\\ 0.27&1&&\\ -0.99&-0.29&1&\\ 0.95&-0.02&-0.95&1\end{pmatrix}$
Figure 1: Boxplots of ratio of means estimations, comparing the MC/MC, CV/MC, and CV/CV estimators
Strong variance reduction with all estimators
As shown in Figure 1, the optimal coefficients $(\alpha_{o},\beta_{o})$ achieve greater variance reduction than alternative choices. This highlights the advantage of jointly optimizing both coefficients for the ratio, rather than optimizing them separately without accounting for the ratio as a whole. The CV/MC estimator does not outperform the CV/CV estimator in terms of variance reduction, further confirming the benefit of optimizing both coefficients simultaneously. Since the CV/CV estimator consistently yielded superior results across experiments, we do not consider the CV/MC estimator in the rest of this work.
4.2 Variance increase with non-optimal coefficients
The presented results are the worst-case scenario for the coefficients of [Gordon et al., 1982], happening with the covariance matrix
$\Sigma=\begin{pmatrix}1&&&\\ -0.99&1&&\\ 0.99&-0.99&1&\\ -0.02&0.02&-0.01&1\end{pmatrix}$
Figure 2: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators
Variance increase with non-optimal coefficients
While the variance reduction achieved with the optimal coefficients may not always exceed that of the other coefficients, their key advantage lies in their theoretical guarantee of variance reduction. The classical coefficients and the ones from [Gordon et al., 1982] can lead to a variance increase as illustrated in Figure 2. This result highlights the importance of jointly optimizing the coefficients to minimize the variance of the whole estimator.
4.3 Best variance reduction with the coefficients of [Gordon et al., 1982]
The presented results are the best-case scenario for the coefficients of [Gordon et al., 1982], with the following covariance matrix
$\Sigma=\begin{pmatrix}1&&&\\ -0.58&1&&\\ -0.99&0.57&1&\\ 0.99&-0.49&-0.99&1\end{pmatrix}$
Figure 3: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators
Best variance reduction with the coefficients from [Gordon et al., 1982] with $n=10$
The optimal coefficients introduced in this work come with a theoretical guarantee of variance reduction and can outperform previously proposed coefficients. However, their complex expressions make them hard to estimate with few samples, as illustrated here with $n=10$ in Figure 3: the optimal coefficients are outperformed by the other ones. Nevertheless, the differences in variance reduction are small and the optimal coefficients give the best results with the same covariance structure and more samples, as seen in Figure 4 with $n=100$.
Figure 4: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators
Equivalent variance reduction performances with $n=100$
5 Application on a multifidelity aircraft design use case
Figure 5: ONERA ALBATROS configuration with strut-braced wings (Figure 11 from [Carrier et al.,])
Figure 6: Illustrations of the $1,252$ HF and LF samples of the dataset
The CV/CV estimator is applied to estimate the strut mass fraction, defined as the ratio between the strut mass and the total mass. As illustrated in Figure 5, the strut is an external support between the fuselage and the wing that helps carry aerodynamic loads. The strut mass fraction is a quantity of interest to be optimized as it ensures the added support does not negate the aerodynamic and structural advantages of a strut-braced wing. The studied dataset contains 1,252 samples from an aircraft design study, based on [Méheut et al., 2024] which describes the multidisciplinary and multi-fidelity design process. The dataset combines outputs from high fidelity (HF) [Lannoo and Priasso, 2025] and low fidelity (LF) [Carrier et al.,] models to efficiently explore the design space. By applying the proposed CV/CV method, the strut mass fraction can be estimated more precisely with fewer HF model evaluations, which take around 2 minutes each, leveraging the LF model, which runs in only a few seconds and serves as a control variate as in MFMC.
Following the notation previously used in this paper, the strut mass fraction can be written as the ratio $R=\mathbb{E}[A]/\mathbb{E}[C]$ where $A$ denotes the HF strut mass and $C$ the HF total mass. The control variates $B$ and $D$ correspond to the LF strut mass and the LF total mass. Since their exact expectations are unknown, we apply the approximate control variates method, estimating their means using the full set of $n+m=1,252$ available samples. Two scenarios are studied: $n=200$ and $n=500$ HF samples. The results are obtained by bootstrapping on 1,000 random dataset configurations.
Simulations highlighted the central role of the correlation structure between the variables and the control variates. The estimated linear correlation matrix for the aircraft design dataset is:
$\begin{pmatrix}1&&&\\ 0.51&1&&\\ 0.77&0.4&1&\\ 0.83&0.78&0.74&1\end{pmatrix}$
(a) $n=200$ HF samples
(b) $n=500$ HF samples
Figure 7: Boxplots of strut mass fraction estimations
As seen in Figure 7, the variance is increased when using the non-optimal coefficients, which highlights the benefit of the proposed method. In both scenarios, our approach achieves an RVR of about 20%, confirming its effectiveness in improving estimation accuracy and delivering the precision of a larger HF dataset at no extra computational cost.
6 Conclusion
In this paper, we introduced new optimized control variates coefficients for a variance-reduced ratio of means estimator. The resulting CV/CV estimator guarantees variance reduction when the coefficients $\alpha$ and $\beta$ are jointly optimized to minimize the overall variance. Significant variance reduction can be achieved, depending on the correlation structure between the variables of interest and the control variates. In particular, we showed that previously proposed coefficients could lead to a variance increase. To illustrate the approach, we applied the estimators with approximate control variates to estimate the strut mass fraction on a multi-fidelity aircraft design dataset.
Future work could aim to apply this method to reduce the variance of conditional mean estimators, such as the tail dependence [Malevergne and Sornette, 2002] or the conditional value at risk [Hong et al., 2014], as well as the self-normalized importance sampling estimator [Bugallo et al., 2017]. The method could also be generalized to estimators defined as functions of ratios. However, the current coefficients do not guarantee variance reduction in that setting; new coefficients would need to be derived for the full estimator.
Supplementary Materials
All simulation codes used to produce the figures are available on GitHub at https://github.com/LouB-N/Control-variates-for-variance-reduced-ratio-of-means-estimators.git.
Acknowledgements
We gratefully acknowledge Eric Nguyen Van (ONERA/DTIS) for providing the dataset used in the application and for his valuable explanations of the aircraft design considerations behind it.
Appendix A Deriving the optimal parameters and proof of variance reduction
To derive the optimal parameters and the variance reduction obtained with the estimators studied, an expression of the variance of a ratio is needed. Denote $X/Z$ the studied ratio, where $X$ and $Z$ are real random variables with non-zero means. Using first-order approximations and the delta method [Casella and Berger, 2024, example 5.5.27] [Sukhatme et al., 1970, section 4a.7] [Cochran, 1977, section 6.9], the variance of the Monte-Carlo estimator $\frac{\overline{X_{n}}}{\overline{Z_{n}}}$ of this ratio can be approximated by:
$\boxed{\text{Var}\left(\frac{\overline{X_{n}}}{\overline{Z_{n}}}\right)=\frac{1}{n}\left(\frac{1}{\mathbb{E}[Z]^{2}}\text{Var}(X)+\frac{\mathbb{E}[X]^{2}}{\mathbb{E}[Z]^{4}}\text{Var}(Z)-2\frac{\mathbb{E}[X]}{\mathbb{E}[Z]^{3}}\text{Cov}\left(X,Z\right)\right)+\mathcal{O}(n^{-2})}$ (31)
Following the same arguments, it can be shown that the approximation for the variance of the CV/MC and CV/CV estimators is of the same order.
A.1 Variance of the MC/MC estimator
An expression of the variance of the MC/MC estimator of a ratio is needed to compare it to the variance of the new variance-reduced estimators.
Recall that the ratio to estimate is
$R=\frac{\mathbb{E}[A]}{\mathbb{E}[C]}$ (32)
and its MC/MC estimator is defined as
$\widehat{R}_{\frac{MC}{MC}}=\frac{\overline{A_{n}}}{\overline{C_{n}}}\qquad\text{ where }\overline{X_{n}}=\frac{1}{n}\sum_{i=1}^{n}X_{i}\quad\text{ for }X\in\{A,C\}$ (33)
To derive the MC/MC estimator’s variance expression, a few preliminary results are needed:
Lemma 1.
  1. 1.
    $\mathbb{E}[\overline{A_{n}}]=\mathbb{E}[A]$;
  2. 2.
    $\mathbb{E}[\overline{C_{n}}]=\mathbb{E}[C]$;
  3. 3.
    $\text{Var}(\overline{A_{n}})=\frac{1}{n}\text{Var}(A)$;
  4. 4.
    $\text{Var}(\overline{C_{n}})=\frac{1}{n}\text{Var}(C)$;
  5. 5.
    $\text{Cov}(\overline{A_{n}},\overline{C_{n}})=\frac{1}{n}\text{Cov}(A,C)$.
Proof.
Points 1 to 4 are standard properties of expectation and variance of averages of i.i.d. random variables.
Point 5 can be proved as
$\text{Cov}(\overline{A_{n}},\overline{C_{n}})=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\text{Cov}(A_{i},C_{j})=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(\text{Cov}(A_{i},C_{i})+\sum_{j\neq i}\text{Cov}(A_{i},C_{j})\right)=\frac{1}{n}\text{Cov}(A,C)$
Proposition 6 (Variance of the MC/MC estimator).
Let $(A_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$ and $C$. The variance of the MC/MC estimator can be approximated as
$\boxed{\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\approx\frac{1}{n\mathbb{E}[C]^{2}}\left(\text{Var}(A)+R^{2}\text{Var}(C)-2RCov(A,C)\right)}$ (34)
Proof.
Using the expressions in Lemma 1, the variance of the MC/MC estimator can be approximated as:
$\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$ $\approx\frac{1}{\mathbb{E}[\overline{C_{n}}]^{2}}\text{Var}(\overline{A_{n}})+\frac{\mathbb{E}[\overline{A_{n}}]^{2}}{\mathbb{E}[\overline{C_{n}}]^{4}}\text{Var}(\overline{C_{n}})-2\frac{\mathbb{E}[\overline{A_{n}}]}{\mathbb{E}[\overline{C_{n}}]^{3}}\text{Cov}(\overline{A_{n}},\overline{C_{n}})$
$\approx\frac{1}{\mathbb{E}[C]^{2}}\text{Var}(\overline{A_{n}})+\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{C_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{C_{n}})$
$\approx\frac{1}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Var}(A)+\frac{R^{2}}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Var}(C)-2\frac{R}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Cov}(A,C)$
A.2 Studying the CV/MC estimator
Recall that the CV/MC estimator is defined as
$\widehat{R}_{\frac{CV}{MC}}=\frac{\widehat{A}_{CV}}{\overline{C_{n}}}=\frac{\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})}{\overline{C_{n}}}$ (35)
A.2.1 Variance of the CV/MC estimator
An expression of the CV/MC estimator’s variance is derived, so that the optimal coefficient minimizing it can be found.
To derive the CV/MC estimator’s variance expression, a few preliminary results are needed:
Lemma 2.
  1. 1.
    $\mathbb{E}[\hat{A}_{CV}]=\mathbb{E}[A]$;
  2. 2.
    $\text{Var}(\hat{A}_{CV})=\frac{1}{n}\text{Var}(A)+\frac{\alpha^{2}}{n}\text{Var}(B)-2\frac{\alpha}{n}\text{Cov}(A,B)$;
  3. 3.
    $\text{Cov}(\hat{A}_{CV},\overline{C_{n}})=\frac{1}{n}\text{Cov}(A,C)-\frac{\alpha}{n}\text{Cov}(B,C)$.
Proof.
Point 1 can be proven as
$\mathbb{E}[\hat{A}_{CV}]=\mathbb{E}[\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})]=\mathbb{E}[\overline{A_{n}}]+\alpha(\mathbb{E}[B]-\mathbb{E}[\overline{B_{n}}])=\mathbb{E}[A]$
Point 2 can be proven as
$\text{Var}(\hat{A}_{CV})$ $=\text{Var}(\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}}))$
$=\text{Var}(\overline{A_{n}})+\alpha^{2}\text{Var}(\mathbb{E}[B]-\overline{B_{n}})+2\alpha\text{Cov}(\overline{A_{n}},\mathbb{E}[B]-\overline{B_{n}})$
$=\text{Var}(\overline{A_{n}})+\alpha^{2}\text{Var}(\overline{B_{n}})-2\alpha\text{Cov}(\overline{A_{n}},\overline{B_{n}})$
$=\frac{1}{n}\text{Var}(A)+\frac{\alpha^{2}}{n}\text{Var}(B)-2\frac{\alpha}{n}\text{Cov}(A,B)$
Point 3 can be proven as
$\text{Cov}(\hat{A}_{CV},\overline{C_{n}})$ $=\text{Cov}(\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}}),\overline{C_{n}})$
$=\text{Cov}(\overline{A_{n}},\overline{C_{n}})+\alpha\text{Cov}(\mathbb{E}[B]-\overline{B_{n}},\overline{C_{n}})$
$=\text{Cov}(\overline{A_{n}},\overline{C_{n}})-\alpha\text{Cov}(\overline{B_{n}},\overline{C_{n}})$
$=\frac{1}{n}\text{Cov}(A,C)-\frac{\alpha}{n}\text{Cov}(B,C)$
Proposition 7 (Variance of the CV/MC estimator).
Let $(A_{i},B_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$, $B$, and $C$, where $B$ is a control variate. The variance of the CV/MC estimator can be approximated as
$\boxed{\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)\approx\frac{1}{n\mathbb{E}[C]^{2}}\left(\text{Var}(A)+\alpha^{2}\text{Var}(B)-2\alpha\text{Cov}(A,B)+R^{2}\text{Var}(C)-2R\text{Cov}(A,C)+2\alpha R\text{Cov}(B,C)\right)}$ (36)
Proof.
Using the expressions in Lemma 2, the variance of the CV/MC estimator can be approximated as:
$\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)$ $\approx\frac{1}{\mathbb{E}[\overline{C_{n}}]^{2}}\text{Var}(\hat{A}_{CV})+\frac{\mathbb{E}[\hat{A}_{CV}]^{2}}{\mathbb{E}[\overline{C_{n}}]^{4}}\text{Var}(\overline{C_{n}})-2\frac{\mathbb{E}[\hat{A}_{CV}]}{\mathbb{E}[\overline{C_{n}}]^{3}}\text{Cov}(\hat{A}_{CV},\overline{C_{n}})$
$\approx\frac{1}{\mathbb{E}[C]^{2}}\text{Var}(\hat{A}_{CV})+\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{C_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\hat{A}_{CV},\overline{C_{n}})$
$\approx\frac{1}{\mathbb{E}[C]^{2}}\text{Var}(\overline{A_{n}})+\frac{\alpha^{2}}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})-\frac{2\alpha}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})+\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{C_{n}})$
$\quad-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{C_{n}})+2\alpha\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})$
$\approx\frac{1}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Var}(A)+\frac{\alpha^{2}}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Var}(B)-\frac{2\alpha}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Cov}(A,B)+\frac{R^{2}}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Var}(C)$
$\quad-2\frac{R}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Cov}(A,C)+2\alpha\frac{R}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Cov}(B,C)$
A.2.2 Computing the optimal coefficient of the CV/MC estimator
Proof of Definition 5.
The goal is to find the optimal coefficient $\alpha_{o}^{\prime}$ minimizing the variance of the CV/MC estimator.
$\alpha_{o}^{\prime}:=\underset{\alpha\in\mathbb{R}}{\text{argmin}}\;\text{Var}\left(\frac{\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})}{\overline{C_{n}}}\right)$
The variance in (36) is minimized by differentiating its expression with respect to $\alpha$ and solving for the optimal value:
$\frac{d}{d\alpha}\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)\approx\frac{2\alpha}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})-\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})+2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})$
The function $\alpha\mapsto\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)$ is strictly convex as $\frac{d^{2}}{d\alpha^{2}}\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)=\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})>0$, so the value of $\alpha$ for which $\frac{d}{d\alpha}\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)=0$ is the unique minimizer of the variance.
It gives:
$\frac{d}{d\alpha}\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)=0\Leftrightarrow\alpha_{o}^{\prime}$ $=\frac{\mathbb{E}[C]^{2}}{2\text{Var}(\overline{B_{n}})}\left(\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})\right)$
$=\frac{\text{Cov}(\overline{A_{n}},\overline{B_{n}})}{\text{Var}(\overline{B_{n}})}-\frac{\mathbb{E}[A]\text{Cov}(\overline{B_{n}},\overline{C_{n}})}{\mathbb{E}[C]\text{Var}(\overline{B_{n}})}$
$\boxed{\alpha_{o}^{\prime}=\frac{\text{Cov}(A,B)}{\text{Var}(B)}-R\frac{\text{Cov}(B,C)}{\text{Var}(B)}}$
A.2.3 Comparing the variances of the MC/MC and the CV/MC estimators with different coefficients
The difference between variances can be expressed as:
$\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$
$\approx$ $\frac{1}{n\mathbb{E}[C]^{2}}\left(\text{Var}(A)+\alpha^{2}\text{Var}(B)-2\alpha\text{Cov}(A,B)+R^{2}\text{Var}(C)-2R\text{Cov}(A,C)+2\alpha R\text{Cov}(B,C)\right)$
$-\frac{1}{n\mathbb{E}[C]^{2}}\left(\text{Var}(A)+R^{2}\text{Var}(C)-2RCov(A,C)\right)$
$\approx$ $\frac{1}{n\mathbb{E}[C]^{2}}\left(\alpha^{2}\text{Var}(B)-2\alpha\text{Cov}(A,B)-2R\text{Cov}(A,C)+2\alpha R\text{Cov}(B,C)\right)$
$\bullet$ With $\alpha_{c}=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ the classical coefficient, the variance difference is
$\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\approx\frac{1}{n\mathbb{E}[C]^{2}}\frac{\text{Cov}(A,B)}{\text{Var}(B)}\left(2R\text{Cov}(B,C)-\text{Cov}(A,B)\right)$
So there is variance reduction if and only if $\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)<0\Leftrightarrow\text{Cov}(A,B)\left(2R\text{Cov}(B,C)-\text{Cov}(A,B)\right)<0$.
$\bullet$ With $\alpha_{o}^{\prime}=\frac{\text{Cov}(A,B)}{\text{Var}(B)}-R\frac{\text{Cov}(B,C)}{\text{Var}(B)}$ the optimized coefficient, the variance difference is
$\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\approx-\frac{1}{n\mathbb{E}[C]^{2}}\frac{\left(\text{Cov}(A,B)-R\text{Cov}(B,C)\right)^{2}}{\text{Var}(B)}\leq 0$
The variance reduction is always guaranteed.
A.3 Studying the CV/CV estimator
Recall that the CV/CV estimator is defined as
$\widehat{R}_{\frac{CV}{CV}}=\frac{\widehat{A}_{CV}}{\overline{C_{CV}}}=\frac{\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})}{\overline{C_{n}}+\alpha(\mathbb{E}[D]-\overline{D_{n}})}$ (37)
A.3.1 Variance of the CV/CV estimator
An expression of the CV/CV estimator’s variance is derived, so that the optimal coefficients minimizing it can be found.
To derive the CV/CV estimator’s variance expression, a few preliminary results are needed:
Lemma 3.
  1. 1.
    $\text{Var}(\hat{A}_{CV})=\text{Var}(\overline{A_{n}})+\alpha^{2}\text{Var}(\overline{B_{n}})-2\alpha\text{Cov}(\overline{A_{n}},\overline{B_{n}})$;
  2. 2.
    $\text{Var}(\hat{C}_{CV,n})=\text{Var}(\overline{C_{n}})+\beta^{2}\text{Var}(\overline{D_{n}})-2\beta\text{Cov}(\overline{C_{n}},\overline{D_{n}})$;
  3. 3.
    $\text{Cov}(\hat{A}_{CV},\hat{C}_{CV,n})=\text{Cov}(\overline{A_{n}},\overline{C_{n}})-\alpha\text{Cov}(\overline{B_{n}},\overline{C_{n}})-\beta\text{Cov}(\overline{A_{n}},\overline{D_{n}})+\alpha\beta\text{Cov}(\overline{B_{n}},\overline{D_{n}})$
Proof.
Points 1 to 3 are standard properties of variance and covariance. ∎
Proposition 8 (Variance of the CV/CV estimator).
Let $(A_{i},B_{i},C_{i},D_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$, $B$, $C$, and $D$, where $B$ and $D$ are control variates. The variance of the CV/CV estimator can be approximated as
$\boxed{\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)\approx\frac{1}{n\mathbb{E}[C]^{2}}\left.\begin{aligned} \Big(&\text{Var}(A)+\alpha^{2}\text{Var}(B)-2\alpha\text{Cov}(A,B)+R^{2}\text{Var}(C)+\beta^{2}R^{2}\text{Var}(D)-2\beta R^{2}\text{Cov}(C,D)\\ &-2R\text{Cov}(A,C)+2\alpha R\text{Cov}(B,C)+2\beta R\text{Cov}(A,D)-2\alpha\beta R\text{Cov}(B,D)\Big)\end{aligned}\right.}$ (38)
Proof.
Using the results from Lemma 3, the CV/CV estimator’s variance is derived as:
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$ $\approx\frac{1}{\mathbb{E}[\hat{C}_{CV,n}]^{2}}\text{Var}(\hat{A}_{CV})+\frac{\mathbb{E}[\hat{A}_{CV}]^{2}}{\mathbb{E}[\hat{C}_{CV,n}]^{4}}\text{Var}(\hat{C}_{CV,n})-2\frac{\mathbb{E}[\hat{A}_{CV}]}{\mathbb{E}[\hat{C}_{CV,n}]^{3}}\text{Cov}\left(\hat{A}_{CV},\hat{C}_{CV,n}\right)$
$\approx\frac{1}{\mathbb{E}[C]^{2}}\text{Var}(\hat{A}_{CV})+\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\hat{C}_{CV,n})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}\left(\hat{A}_{CV},\hat{C}_{CV,n}\right)$
$\approx\frac{1}{\mathbb{E}[C]^{2}}\text{Var}(\overline{A_{n}})+\frac{\alpha^{2}}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})-\frac{2\alpha}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})$
$\quad+\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{C_{n}})+\beta^{2}\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})-2\beta\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(\overline{C_{n}},\overline{D_{n}})$
$\quad-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{C_{n}})+2\alpha\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})+2\beta\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{D_{n}})-2\alpha\beta\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})$
$\approx\frac{1}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Var}(A)+\frac{\alpha^{2}}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Var}(B)-\frac{2\alpha}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Cov}(A,B)$
$\quad+\frac{R^{2}}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Var}(C)+\beta^{2}\frac{R^{2}}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Var}(D)-2\beta\frac{R^{2}}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Cov}(C,D)$
$\quad-2\frac{R}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Cov}(A,C)+2\alpha\frac{R}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Cov}(B,C)+2\beta\frac{R}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Cov}(A,D)-2\alpha\beta\frac{R}{\mathbb{E}[C]^{2}}\frac{1}{n}\text{Cov}(B,D)$
A.3.2 Computing the optimal coefficients of the CV/CV estimator
Proof of Proposition 2.
$(\alpha_{o},\beta_{o}):=\underset{(\alpha,\beta)\in\mathbb{R}^{2}}{\text{argmin}}\;\text{Var}\left(\frac{\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})}{\overline{C_{n}}+\beta(\mathbb{E}[D]-\overline{D_{n}})}\right)$
The variance in (38) is minimized by differentiating its expression with respect to $\alpha$ and $\beta$, and solving for the optimal value:
$\nabla\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$ $=\begin{bmatrix}\frac{\partial}{\partial\alpha}\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)\\ \frac{\partial}{\partial\beta}\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)\end{bmatrix}$
$=\begin{bmatrix}\frac{2\alpha}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})-\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})+2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})-2\beta\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\\ 2\beta\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})-2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(\overline{C_{n}},\overline{D_{n}})+2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{D_{n}})-2\alpha\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\end{bmatrix}$
The function $(\alpha,\beta)\mapsto\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$ is strictly convex as
$\nabla^{2}\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$ $=\begin{bmatrix}\frac{\partial^{2}}{\partial\alpha^{2}}\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)&\frac{\partial^{2}}{\partial\alpha\partial\beta}\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)\\ \frac{\partial^{2}}{\partial\beta\partial\alpha}\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)&\frac{\partial^{2}}{\partial\beta^{2}}\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)\end{bmatrix}$
$=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})&-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\\ -2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})&2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})\end{bmatrix}:=M$
and M is positive definite as
$det(M)>0$
$\Leftrightarrow$ $\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})\times 2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})-\left(-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\right)^{2}>0$
$\Leftrightarrow$ $\frac{4\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{6}}\text{Var}(\overline{B_{n}})\text{Var}(\overline{D_{n}})-\frac{4\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{6}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})^{2}>0$
$\Leftrightarrow$ $\text{Var}(\overline{B_{n}})\text{Var}(\overline{D_{n}})>\text{Cov}(\overline{B_{n}},\overline{D_{n}})^{2}$
$\Leftrightarrow$ $|\text{Corr}(B,D)|<1$
So the values of $\alpha_{o}$ and $\beta_{o}$ for which $\nabla\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)=\begin{bmatrix}0\\ 0\end{bmatrix}$ is the unique minimizer of the variance, under the mild condition $|\text{Corr}(B,D)|<1$. Other settings with $|\text{Corr}(B,D)|=1$ are explored in the next section.
To find $\alpha_{o}$ and $\beta_{o}$, the following system need to be solved:
$\nabla\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)=\begin{bmatrix}0\\ 0\end{bmatrix}\Leftrightarrow M\cdot\begin{bmatrix}\alpha_{o}\\ \beta_{o}\end{bmatrix}=\begin{bmatrix}c_{1}\\ c_{2}\end{bmatrix}$
with $M=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})&-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\\ -2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})&2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})\end{bmatrix}$ and $\begin{bmatrix}c_{1}\\ c_{2}\end{bmatrix}=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})\\ 2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(\overline{C_{n}},\overline{D_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{D_{n}})\end{bmatrix}$.
The solution is given by: $\begin{bmatrix}\alpha_{o}\\ \beta_{o}\end{bmatrix}=M^{-1}\cdot\begin{bmatrix}c_{1}\\ c_{2}\end{bmatrix}$, where the inverse of $M$ is:
$M^{-1}=\frac{1}{\frac{4\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{6}}\text{Var}(\overline{B_{n}})\text{Var}(\overline{D_{n}})-\frac{4\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{6}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})^{2}}\begin{bmatrix}2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})&2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\\ 2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})&\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})\end{bmatrix}$.
The coefficients can be derived as follows:
$\alpha_{o}$ $=\frac{2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})\left(\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})\right)+2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\left(2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(\overline{C_{n}},\overline{D_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{D_{n}})\right)}{\frac{4\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{6}}\text{Var}(\overline{B_{n}})\text{Var}(\overline{D_{n}})-\frac{4\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{6}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})^{2}}$
$=\frac{2\frac{R^{2}}{\mathbb{E}[C]^{2}}\text{Var}(D)\left(\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(A,B)-2\frac{R}{\mathbb{E}[C]^{2}}\text{Cov}(B,C)\right)+2\frac{R}{\mathbb{E}[C]^{2}}\text{Cov}(B,D)\left(2\frac{R^{2}}{\mathbb{E}[C]^{2}}\text{Cov}(C,D)-2\frac{R}{\mathbb{E}[C]^{2}}\text{Cov}(A,D)\right)}{\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Var}(B)\text{Var}(D)-\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(B,D)^{2}}$
$=\frac{\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Var}(D)\text{Cov}(A,B)-\frac{4R^{3}}{\mathbb{E}[C]^{4}}\text{Var}(D)\text{Cov}(B,C)+\frac{4R^{3}}{\mathbb{E}[C]^{4}}\text{Cov}(B,D)\text{Cov}(C,D)-\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(B,D)\text{Cov}(A,D)}{\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Var}(B)\text{Var}(D)-\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(B,D)^{2}}$
$=\frac{\text{Var}(D)\text{Cov}(A,B)-R\text{Var}(D)\text{Cov}(B,C)+R\text{Cov}(B,D)\text{Cov}(C,D)-\text{Cov}(B,D)\text{Cov}(A,D)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}$
$\beta_{o}$ $=\frac{2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\left(\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})\right)+\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})\left(2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(\overline{C_{n}},\overline{D_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{D_{n}})\right)}{\frac{4\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{6}}\text{Var}(\overline{B_{n}})\text{Var}(\overline{D_{n}})-\frac{4\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{6}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})^{2}}$
$=\frac{2\frac{R}{\mathbb{E}[C]^{2}}\text{Cov}(B,D)\left(\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(A,B)-2\frac{R}{\mathbb{E}[C]^{2}}\text{Cov}(B,C)\right)+\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(B)\left(2\frac{R^{2}}{\mathbb{E}[C]^{2}}\text{Cov}(C,D)-2\frac{R}{\mathbb{E}[C]^{2}}\text{Cov}(A,D)\right)}{\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Var}(B)\text{Var}(D)-\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(B,D)^{2}}$
$=\frac{\frac{4R}{\mathbb{E}[C]^{4}}\text{Cov}(B,D)\text{Cov}(A,B)-\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(B,D)\text{Cov}(B,C)+\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Var}(B)\text{Cov}(C,D)-\frac{4R}{\mathbb{E}[C]^{4}}\text{Var}(B)\text{Cov}(A,D)}{\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Var}(B)\text{Var}(D)-\frac{4R^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(B,D)^{2}}$
$=\frac{\frac{1}{R}\text{Cov}(B,D)\text{Cov}(A,B)-\text{Cov}(B,D)\text{Cov}(B,C)+\text{Var}(B)\text{Cov}(C,D)-\frac{1}{R}\text{Var}(B)\text{Cov}(A,D)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}$
The optimal coefficients are:
$\boxed{\alpha_{o}=\frac{\text{Var}(D)\text{Cov}(A,B)-R\text{Var}(D)\text{Cov}(B,C)+R\text{Cov}(B,D)\text{Cov}(C,D)-\text{Cov}(B,D)\text{Cov}(A,D)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}}$ (39)
$\boxed{\beta_{o}=\frac{\frac{1}{R}\text{Cov}(B,D)\text{Cov}(A,B)-\text{Cov}(B,D)\text{Cov}(B,C)+\text{Var}(B)\text{Cov}(C,D)-\frac{1}{R}\text{Var}(B)\text{Cov}(A,D)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}}$ (40)
A.3.3 Computing the optimal coefficients of the CV/CV estimator if the control variates $B$ and $D$ are linearly correlated
Proof of Proposition 4.
The setting where $\text{Corr}(B,D)=1$ is studied in this section. The two control variates $B$ and $D$ are linearly correlated if there is a linear relationship between them, i.e. $B=aD+b$ where $(a,b)\in\mathbb{R}^{*}\times\mathbb{R}$.
Some terms in the variance expression involving $B$ can be written in terms of $D$:
$\mathbb{E}[B]$ $=a\mathbb{E}[D]+b$
$\text{Var}(B)$ $=a^{2}\text{Var}(D)$
$\text{Cov}(X,B)$ $=a\text{Cov}(X,D)\quad\forall X\in\{A,C\}$
The CV/CV estimator’s variance can be expressed as:
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)\approx\frac{1}{n\mathbb{E}[C]^{2}}\left.\begin{aligned} \Big(&\text{Var}(A)+a^{2}\alpha^{2}\text{Var}(D)-2a\alpha\text{Cov}(A,D)+R^{2}\text{Var}(C)+\beta^{2}R^{2}\text{Var}(D)-2\beta R^{2}\text{Cov}(C,D)\\ &-2R\text{Cov}(A,C)+2a\alpha R\text{Cov}(C,D)+2\beta R\text{Cov}(A,D)-2a\alpha\beta R\text{Var}(D)\Big)\end{aligned}\right$.
This variance is minimized by differentiating its expression with respect to $\alpha$ and $\beta$, and solving for the optimal value:
$\nabla\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$ $=\begin{bmatrix}\frac{\partial}{\partial\alpha}\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)\\ \frac{\partial}{\partial\beta}\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)\end{bmatrix}$
$=\frac{2}{n\mathbb{E}[C]^{2}}\begin{bmatrix}a^{2}\alpha\text{Var}(D)-a\text{Cov}(A,D)+aR\text{Cov}(C,D)-a\beta R\text{Var}(D)\\ \beta R^{2}\text{Var}(D)-R^{2}\text{Cov}(C,D)+R\text{Cov}(A,D)-a\alpha R\text{Var}(D)\end{bmatrix}$
$\nabla\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)=\begin{bmatrix}0\\ 0\end{bmatrix}$ is solved to find the expression of the minimizers $\alpha_{o}$ and $\beta_{o}$.
$\nabla\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)=\begin{bmatrix}0\\ 0\end{bmatrix}$ $\Leftrightarrow\begin{cases}\text{Var}(D)\left(a\alpha_{o}-\beta_{o}R\right)=\text{Cov}(A,D)-R\text{Cov}(C,D)\\ \text{Var}(D)\left(\beta_{o}R-a\alpha_{o}\right)=R\text{Cov}(C,D)-\text{Cov}(A,D)\end{cases}$
$\Leftrightarrow\alpha_{o}=\frac{1}{a}\left(\frac{\text{Cov}(A,D)-R\text{Cov}(C,D)}{\text{Var}(D)}+\beta_{o}R\right)$
If the control variates $B$ and $D$ are linearly correlated, there is an infinity of minimizers defined as
$\begin{cases}\alpha_{o}=\frac{1}{a}\left(\frac{\text{Cov}(A,D)-R\text{Cov}(C,D)}{\text{Var}(D)}+\beta_{o}R\right)\\ \beta_{o}\in\mathbb{R}\end{cases}$ (41)
Note $\eta:=\frac{\text{Cov}(A,D)-R\text{Cov}(C,D)}{\text{Var}(D)}$. By replacing $\alpha$ and $\beta$ with their optimal values $\alpha_{o}$ and $\beta_{o}$, the variance expression is as follows:
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$ $\approx\frac{1}{n\mathbb{E}[C]^{2}}\left.\begin{aligned} \Big(&\text{Var}(A)+\left(\eta+\beta_{o}R\right)^{2}\text{Var}(D)-2\left(\eta+\beta_{o}R\right)\text{Cov}(A,D)+R^{2}\text{Var}(C)+\beta_{o}^{2}R^{2}\text{Var}(D)\\ &-2\beta_{o}R^{2}\text{Cov}(C,D)-2R\text{Cov}(A,C)+2\left(\eta+\beta_{o}R\right)R\text{Cov}(C,D)+2\beta_{o}R\text{Cov}(A,D)\\ &-2\left(\eta+\beta_{o}R\right)\beta_{o}R\text{Var}(D)\Big)\end{aligned}\right$.
$\approx\frac{1}{n\mathbb{E}[C]^{2}}\left.\begin{aligned} \Big(&\text{Var}(A)+\eta^{2}\text{Var}(D)-2\eta\text{Cov}(A,D)+R^{2}\text{Var}(C)-2R\text{Cov}(A,C)+2\eta R\text{Cov}(C,D)\\ &+2\eta\beta_{o}R\text{Var}(D)+\beta_{o}^{2}R^{2}\text{Var}(D)-2\beta_{o}R\text{Cov}(A,D)+\beta_{o}^{2}R^{2}\text{Var}(D)-2\beta_{o}R^{2}\text{Cov}(C,D)\\ &+2\beta_{o}R^{2}\text{Cov}(C,D)+2\beta_{o}R\text{Cov}(A,D)-2\eta\beta_{o}R\text{Var}(D)-2\beta_{o}^{2}R^{2}\text{Var}(D)\Big)\end{aligned}\right$.
$\approx\frac{1}{n\mathbb{E}[C]^{2}}\left(\text{Var}(A)+\eta\text{Var}(D)-2\eta\text{Cov}(A,D)+R^{2}\text{Var}(C)-2R\text{Cov}(A,C)+2\eta R\text{Cov}(C,D)\right)$
The minimized variance is the same for any value of $\beta_{o}$. Note that it is independent of $a$ and $b$, which means the reduced variance is the same for any control variates $B$ and $D=aB+d$.
The setting with two equal control variates $B=D$ is included in this study with $a=1$ and $b=0$.
The variance reduction with linearly correlated control variates and optimal coefficients takes the following form:
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$
$\approx\frac{1}{n\mathbb{E}[C]^{2}}\left(\eta\text{Var}(D)-2\eta\text{Cov}(A,D)+2\eta R\text{Cov}(C,D)\right)$
$\approx\frac{\text{Cov}(A,D)\text{Var}(D)-R\text{Cov}(C,D)\text{Var}(D)-2\text{Cov}(A,D)^{2}+4R\text{Cov}(C,D)\text{Cov}(A,D)-2R^{2}\text{Cov}(C,D)^{2}}{n\mathbb{E}[C]^{2}\text{Var}(D)}$
$\approx\frac{1}{n\mathbb{E}[C]^{2}}\left(\text{Cov}(A,D)-R\text{Cov}(C,D)-2\frac{\left(\text{Cov}(A,D)-R\text{Cov}(C,D)\right)^{2}}{\text{Var}(D)}\right)$
$\approx\frac{1}{n\mathbb{E}[C]^{2}}\tau\left(1-2\frac{\tau^{2}}{\text{Var}(D)}\right)\quad\text{where }\tau=\text{Cov}(A,D)-R\text{Cov}(C,D)=\text{Cov}(A-RC,D)$
The variance is reduced if $\tau\left(1-2\frac{\tau^{2}}{\text{Var}(D)}\right)<0$. That is
[…]
$\Leftrightarrow$ $\begin{cases}-\sqrt{\frac{\text{Var}(D)}{2}}<\text{Cov}(A-RC,D)<0\\ \sqrt{\frac{\text{Var}(D)}{2}}<\text{Cov}(A-RC,D)\end{cases}$
Unlike the more general case where the control variates are non linearly correlated, when $B=aD+b$, the variance reduction is not guaranteed unless the condition $\text{Cov}(A-RC,D)\in\left(-\sqrt{\frac{\text{Var}(D)}{2}},0\right)\cup\left(\sqrt{\frac{\text{Var}(D)}{2}},+\infty\right)$ is fulfilled.
A.3.4 Comparing the variances of the MC/MC and the CV/CV estimators with different coefficients
The difference between variances can be expressed as:
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\approx\frac{1}{n\mathbb{E}[C]^{2}}\left(\begin{aligned} &\alpha^{2}\text{Var}(B)-2\alpha\text{Cov}(A,B)+\beta^{2}R^{2}\text{Var}(D)-2\beta R^{2}\text{Cov}(C,D)\\ &\quad+2\alpha R\text{Cov}(B,C)+2\beta R\text{Cov}(A,D)-2\alpha\beta R\text{Cov}(B,D)\end{aligned}\right)$
$\bullet$ With the classical coefficients $\alpha_{c}=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ and $\beta_{c}=\frac{\text{Cov}(C,D)}{\text{Var}(D)}$, the variance difference is:
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\approx$ $\frac{1}{n\mathbb{E}[C]^{2}\text{Var}(B)\text{Var}(D)}\left.\begin{aligned} \big(&-\text{Cov}(A,B)^{2}\text{Var}(D)-R^{2}\text{Cov}(C,D)^{2}\text{Var}(B)\\ &+2R\text{Cov}(A,B)\text{Cov}(B,C)\text{Var}(D)\\ &+2R\text{Cov}(C,D)\text{Cov}(A,D)\text{Var}(B)\\ &-2R\text{Cov}(A,B)\text{Cov}(C,D)\text{Cov}(B,D)\;\big)\end{aligned}\right$.
$:\approx$ $\frac{1}{n\mathbb{E}[C]^{2}}\frac{\textbf{T}}{\text{Var}(B)\text{Var}(D)}$
There is a variance reduction if $\textbf{T}<0$.
$\bullet$ With the coefficients from [Gordon et al., 1982] $\alpha_{g}=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ and $\beta_{g}=\frac{\text{Cov}(C,D)-\frac{1}{R}\text{Cov}(A,D)+\frac{\alpha}{R}\text{Cov}(B,D)}{\text{Var}(D)}$, the variance difference is:
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$
$\approx\frac{1}{n\mathbb{E}[C]^{2}}\bigg(\frac{\textbf{T}-\text{Cov}(A,D)^{2}Var(B)+2Cov(A,B)\text{Cov}(B,D)\text{Cov}(A,D)}{\text{Var}(B)\text{Var}(D)}-\frac{\text{Cov}(A,B)^{2}Cov(B,D)^{2}}{\text{Var}(B)^{2}Var(D)}\bigg)$
There is a variance reduction if $\text{Var}(B)\bigg(\textbf{T}-\text{Cov}(A,D)^{2}Var(B)+2Cov(A,B)\text{Cov}(B,D)\text{Cov}(A,D)\bigg)-\text{Cov}(A,B)^{2}Cov(B,D)^{2}<0$.
$\bullet$ With the optimal coefficients αo=Var(D)Cov(A,B)RVar(D)Cov(B,C)+RCov(B,D)Cov(C,D)Cov(B,D)Cov(A,D)Var(B)Var(D)Cov(B,D)2 and
βo
=1RCov(B,D)Cov(A,B)Cov(B,D)Cov(B,C)+Var(B)Cov(C,D)1RVar(B)Cov(A,D)Var(B)Var(D)Cov(B,D)2
[…]
, the variance difference is:
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$ $\approx-\frac{1}{n\mathbb{E}[C]^{2}}\frac{Var\bigg(\big(R\text{Cov}(B,C)-\text{Cov}(A,B)\big)D-\big(R\text{Cov}(C,D)-\text{Cov}(A,D)\big)B\bigg)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}\leq 0$
The variance reduction is guaranteed.
A.4 Calculs approximate CV
Recall that the exact control variate estimator of $\mathbb{E}[A]$ is $\widehat{A}_{CV}=\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})$ and the approximate estimator is $\widehat{A}_{ACV}=\overline{A_{n}}+\alpha(\overline{B_{n+m}}-\overline{B_{n}})$, where $B$ is the control variate.
Considering the following properties:
$\text{Var}(\overline{B_{n+m}}-\overline{B_{n}})=\frac{m}{n+m}\text{Var}(\mathbb{E}[B]-\overline{B_{n}})\qquad\qquad\quad\text{Var}(\overline{D_{n+m}}-\overline{D_{n}})=\frac{m}{n+m}\text{Var}(\mathbb{E}[D]-\overline{D_{n}})$
$\text{Cov}(\overline{Y_{n+m}}-\overline{Y_{n}},\overline{X_{n}})=\frac{m}{n+m}\text{Cov}(\mathbb{E}[Y]-\overline{Y_{n}},\overline{X_{n}})\qquad X\in\{A,C\},Y\in\{B,D\}$
$\text{Cov}(\overline{B_{n+m}}-\overline{B_{n}},\overline{D_{n+m}}-\overline{D_{n}})=\frac{m}{n+m}\text{Cov}(\mathbb{E}[B]-\overline{B_{n}},\mathbb{E}[D]-\overline{D_{n}})$
The coefficients that minimize the variance remain the same in the approximate case as in the exact case.
It can easily be shown that the variance reductions in the approximate case are the same as in the exact case, with an additional factor of $\frac{m}{n+m}$ i.e.
$\text{Var}\left(\widehat{R}_{\frac{ACV}{MC}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)=\frac{m}{n+m}\left(\text{Var}\left(\widehat{R}_{\frac{CV}{MC}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\right)$ (42)
$\text{Var}\left(\widehat{R}_{\frac{ACV}{ACV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)=\frac{m}{n+m}\left(\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\right)$ (43)
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, the authors used ChatGPT in order to improve language and readability. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.

### v2

@@
 Jérôme Morio ONERA/DTIS, Université de Toulouse, F-31055 Toulouse Fédération ENAC ISAE-SUPAERO ONERA, Université de Toulouse, 31000 Toulouse Benjamin Bobbia Fédération ENAC ISAE-SUPAERO ONERA, Université de Toulouse, 31000 Toulouse
-( Preprint - October 15, 2025)
+( Preprint - November 7, 2025)
 The control variates method is a classical variance reduction technique for Monte Carlo estimators that exploits correlated auxiliary variables without introducing bias.
@@
 Simulation studies show significant variance reduction, particularly when correlations between variables and control variates are strong.
-The practical value of the method is illustrated with a multi-fidelity aircraft design use case.
+The practical value of the method is illustrated on multi-fidelity applications: estimating a proportion in an aircraft design use case and a conditional value-at-risk in an electromagnetic dataset.
 1 Introduction
@@
 As a result, every observation has a large influence on the estimate, leading to high variance and reduced reliability.
-An overview of common variance reduction techniques can be found in [Asmussen and Glynn, 2007].
+An overview of common variance reduction techniques can be found in [asmussen_variance-reduction_2007].
 In the setting considered here, where only a limited number of samples is available and further sampling is not possible, most methods are not directly applicable.
-The control variates method [Asmussen and Glynn, 2007, Sec.
+The control variates method [asmussen_variance-reduction_2007, Sec.
 2] is well-suited, as it exploits i.i.d.
 samples from an auxiliary variable correlated with the target variable, to reduce the variance of the Monte Carlo estimator without introducing bias.
-It can be seen as a form of transfer learning [Zhu et al., 2025], where information from a source variable is used to improve estimation on the target variable.
-One of the first references to the control variates method appears in [Hammersley and Handscomb, 1965], Section 2.6, where it is introduced under the name of “regression estimator.” The method can be interpreted as a linear regression, where the coefficient of the control variate is chosen as the slope that best explains the variability of the output through its correlation with the auxiliary variable.
-Later works developed this regression view further: [Lewis et al., 1987] with linear and nonlinear adjustments, [Davidson and MacKinnon, 1992] with applications to quantile estimation, and [Salaün et al., 2022] with regression-based Monte Carlo integration.
+It can be seen as a form of transfer learning [zhu_recent_2025], where information from a source variable is used to improve estimation on the target variable.
+One of the first references to the control variates method appears in [hammersley_monte_1965], Section 2.6, where it is introduced under the name of “regression estimator.” The method can be interpreted as a linear regression, where the coefficient of the control variate is chosen as the slope that best explains the variability of the output through its correlation with the auxiliary variable.
+Later works developed this regression view further: [lewis_smarter_1987] with linear and nonlinear adjustments, [davidson_regression-based_1992] with applications to quantile estimation, and [salaun_regression-based_2022] with regression-based Monte Carlo integration.
 Control variates have been adapted in many directions, from theoretical analyses to practical implementations; in the following, we mention a selection of approaches.
-[Hesterberg and Nelson, 1998] discusses the choice of control variates when estimating probabilities or quantiles.
-[Glynn and Szechtman, 2002] provides a comprehensive Hilbert-space perspective on control variates, establishes asymptotic optimality results, and explores their connections with other variance reduction techniques.
-[Hickernell et al., 2005] extends the control variates method to quasi-Monte Carlo estimators, highlighting that the optimal control variate differs from the classical Monte Carlo setting.
-[Müller et al., 2020] introduces Neural Control Variates, using a set of neural networks to apply control variates to parametric Monte Carlo integration.
-[Fox and Ökten, 2021] presents a polynomial chaos expansion method for control variates.
-[Maskell et al., 2022] extends the method to the estimation of constrained variables.
+[hesterberg_control_1998] discusses the choice of control variates when estimating probabilities or quantiles.
+[glynn_new_2002] provides a comprehensive Hilbert-space perspective on control variates, establishes asymptotic optimality results, and explores their connections with other variance reduction techniques.
+[hickernell_control_2005] extends the control variates method to quasi-Monte Carlo estimators, highlighting that the optimal control variate differs from the classical Monte Carlo setting.
+[muller_neural_2020] introduces Neural Control Variates, using a set of neural networks to apply control variates to parametric Monte Carlo integration.
+[fox_polynomial_2021] presents a polynomial chaos expansion method for control variates.
+[maskell_control_2022] extends the method to the estimation of constrained variables.
 The approximate control variates method extends the standard control variates approach to cases where the mean of the auxiliary variable is unknown, estimating it from additional samples.
-It is also referred to as Multi-Fidelity Monte Carlo (MFMC) [Cannamela et al., 2008] [Gorodetsky et al., 2020] [Kim et al., 2024]: the output of a computationally cheaper lower-fidelity model is used as a control variate to improve the estimation on the output of a more accurate high-fidelity model.
+It is also referred to as Multi-Fidelity Monte Carlo (MFMC) [cannamela_controlled_2008] [gorodetsky_generalized_2020] [kim_parametric_2024]: the output of a computationally cheaper lower-fidelity model is used as a control variate to improve the estimation on the output of a more accurate high-fidelity model.
 Since both models approximate the same underlying quantity, their outputs are often highly correlated, which enables efficient variance reduction through control variates.
@@
 The standard approach consists of estimating the numerator and denominator separately by Monte Carlo, and then taking their ratio.
-This estimator has been analysed in [Cochran, 1977].
+This estimator has been analysed in [cochran_sampling_1977].
 The control variates approach can be applied to both components to reduce variance.
-To the best of our knowledge, [Gordon et al., 1982] is the only work specifically extending the method to ratio of means estimation, by setting the numerator coefficient to its classical value, the one minimizing variance for a single mean, and optimizing the denominator coefficient to minimize the ratio variance.
-The classical coefficients for mean estimation, or those of [Gordon et al., 1982], can in fact increase the variance of the ratio of means estimator with certain correlation structures, as they are not both derived to minimize the variance of the ratio.
+To the best of our knowledge, [gordon_efficient_1982] is the only work specifically extending the method to ratio of means estimation, by setting the numerator coefficient to its classical value, the one minimizing variance for a single mean, and optimizing the denominator coefficient to minimize the ratio variance.
+The classical coefficients for mean estimation, or those of [gordon_efficient_1982], can in fact increase the variance of the ratio of means estimator with certain correlation structures, as they are not both derived to minimize the variance of the ratio.
 This paper introduces new optimal coefficients that theoretically guarantee variance reduction as they are jointly optimized to minimize the variance of the estimator.
@@
 In Section 4, simulation results are presented, illustrating extreme correlation structures and their impact on variance reduction.
-Finally, in Section 5, the proposed estimators are applied to a multi-fidelity aircraft design dataset.
+Finally, in Section 5.1, we demonstrate the proposed estimators on two multi-fidelity datasets: to estimate a proportion in an aircraft design dataset, and a conditional value-at-risk on an electromagnetic dataset.
 2 Variance reduction with control variates
@@
 The control variates estimator of $\mathbb{E}[A]$ is defined as follows.
-Definition 2 (Control variates estimator [Asmussen and Glynn, 2007]).
+Definition 2 (Control variates estimator [asmussen_variance-reduction_2007]).
 Let $(A_{i},B_{i})_{i=1\dots n}$ be i.i.d.
-real samples from the joint distribution of the random variables $A$ and $B$.
+real samples from the joint distribution of the random variables $A,B\in L^{2}$.
 Assume $\mathbb{E}[B]$ is known.
@@
 The coefficient $\alpha_{c}$ needs to be estimated, which can introduce a small bias if the estimation is made from the same dataset used to compute $\widehat{A}_{CV}$.
-This plug-in bias [Kennedy, 2024] is generally negligible in practice, as discussed in [Owen, 2013].
+This plug-in bias [kennedy2024semiparametric] is generally negligible in practice, as discussed in [owen_variance_2013].
 With the optimal coefficient $\alpha_{c}$, the variance of the control variates estimator is
@@
 The approximate control variates (ACV) method overcomes this by estimating the mean from additional data.
-Definition 3 (Approximate control variates [Gorodetsky et al., 2020]).
+Definition 3 (Approximate control variates [gorodetsky_generalized_2020]).
 Let $(A_{i},B_{i})_{i=1\dots n}$ be i.i.d.
-real samples from the joint distribution of the random variables $A$ and $B$, and let $(B_{i})_{i=n+1\dots n+m}$ be additional i.i.d.
+real samples from the joint distribution of the random variables $A,B\in L^{2}$, and let $(B_{i})_{i=n+1\dots n+m}$ be additional i.i.d.
 samples of $B$.
@@
 […]
-The control variate $B$ is typically chosen to mirror the variable it controls [Hesterberg and Nelson, 1998].
-While the method naturally extends to multiple control variates using linear regression to estimate the various coefficients [Leluc et al., 2021], only a single control variate is considered here for clarity.
+The control variate $B$ is typically chosen to mirror the variable it controls [hesterberg_control_1998].
+While the method naturally extends to multiple control variates using linear regression to estimate the various coefficients [leluc_control_2021], only a single control variate is considered here for clarity.
 2.2 Control variates applied to ratio of means estimators
@@
 Let $(A_{i},B_{i},C_{i},D_{i})_{i=1\dots n}$ be i.i.d.
-real samples from the joint distribution of the random variables $A$, $B$, $C$, and $D$, where $B$ and $D$ are control variates.
+real samples from the joint distribution of the random variables $A,B,C,D\in L^{2}$, where $B$ and $D$ are control variates.
 The CV/CV estimator of $R$ is defined as
@@
     $\beta_{c}:=\underset{\beta\in\mathbb{R}}{\text{argmin}}\;\text{Var}\left(\overline{C_{n}}+\beta(\mathbb{E}[D]-\overline{D_{n}})\right)=\frac{\text{Cov}(C,D)}{\text{Var}(D)}$ (12)
-  • the coefficients introduced in [Gordon et al., 1982]:
+  • the coefficients introduced in [gordon_efficient_1982]:
     $\alpha_{g}$ $:=\underset{\alpha\in\mathbb{R}}{\text{argmin}}\;\text{Var}\left(\overline{A_{n}}+\alpha(\mathbb{E}[B]-\overline{B_{n}})\right)$ $=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ (13)
@@
 However, these choices are not always optimal: variance reduction is not guaranteed in all cases, and the variance can even be increased as illustrated analytically in A.3 and with simulations in Section 4.
-The variance difference between the CV/CV and MC/MC estimators can be explicitly expressed, but analysing the conditions on the covariances and variances of $A$, $B$, $C$, and $D$ that ensure variance reduction remains a delicate task.
+The variance difference between the CV/CV and MC/MC estimators can be explicitly expressed, but analyzing the conditions on the covariances and variances of $A$, $B$, $C$, and $D$ that ensure variance reduction remains a delicate task.
 3 Optimized coefficients for control variates ratio of means estimators
@@
 With the optimal coefficients given in (15) and (16), the variance reduction obtained with the CV/CV estimator is:
-$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$ $\approx-\frac{1}{n\mathbb{E}[C]^{2}}\frac{Var\bigg(\big(R\text{Cov}(B,C)-\text{Cov}(A,B)\big)D-\big(R\text{Cov}(C,D)-\text{Cov}(A,D)\big)B\bigg)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}\leq 0$ (17)
+$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$
+$\approx-\frac{1}{n\mathbb{E}[C]^{2}}\frac{Var\bigg(\big(R\text{Cov}(B,C)-\text{Cov}(A,B)\big)D-\big(R\text{Cov}(C,D)-\text{Cov}(A,D)\big)B\bigg)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}\leq 0$ (17)
 The variance reduction is guaranteed:
@@
 Let $(A_{i},B_{i},C_{i})_{i=1\dots n}$ be i.i.d.
-real samples from the joint distribution of the random variables $A$, $B$, and $C$, and let $(B_{i})_{i=n+1\dots n+m}$ be additional i.i.d.
+real samples from the joint distribution of the random variables $A,B,C\in L^{2}$, and let $(B_{i})_{i=n+1\dots n+m}$ be additional i.i.d.
 samples of $B$.
@@
 Simulation results in Section 4.1 confirm that the CV/CV estimator consistently yields lower variance than the CV/MC estimator.
-The MC/MC estimator $\widehat{R}_{\frac{MC}{MC}}$ is asymptotically unbiased as $n\to\infty$ [Cochran, 1977], a property that extends to the control variates estimators $\widehat{R}_{\frac{CV}{MC}}$ and $\widehat{R}_{\frac{CV}{CV}}$ introduced previously.
+The MC/MC estimator $\widehat{R}_{\frac{MC}{MC}}$ is asymptotically unbiased as $n\to\infty$ [cochran_sampling_1977], a property that extends to the control variates estimators $\widehat{R}_{\frac{CV}{MC}}$ and $\widehat{R}_{\frac{CV}{CV}}$ introduced previously.
 For smaller sample sizes, bias may arise due to the nonlinearity of the ratio form: the expectation of a ratio is generally not equal to the ratio of expectations.
 This bias is particularly relevant when the denominator has high variance or is near zero.
-Nonetheless, as illustrated in Sections 4 and 5, the impact of this bias is negligible.
+Nonetheless, as illustrated in Sections 4 and 5.1, the impact of this bias is negligible.
 3.3 Approximate control variates for a ratio of means estimator
@@
   • $(\alpha_{c},\beta_{c})$ the "classical" coefficients, given in (11) and (12), derived by optimizing the numerator and denominator independently;
-  • $(\alpha_{g},\beta_{g})$ the coefficients proposed by [Gordon et al., 1982], given in (13) and (14), where $\alpha_{g}$ minimizes the numerator variance and $\beta_{g}$ minimizes the variance of the ratio;
+  • $(\alpha_{g},\beta_{g})$ the coefficients proposed by [gordon_efficient_1982], given in (13) and (14), where $\alpha_{g}$ minimizes the numerator variance and $\beta_{g}$ minimizes the variance of the ratio;
   • $(\alpha_{o},\beta_{o})$ the new optimal coefficients, given in (15) and (16), that jointly minimize the variance of the full ratio.
@@
 A positive RVR indicates that the new estimator $\widehat{R}_{\text{new}}$ achieves a lower variance than the baseline, while a negative RVR implies a variance increase.
-The random variables $A,B,C,$ and $D$ are modelled as components of a four-dimensional Gaussian vector:
-$(A,B,C,D)^{\top}\sim\mathcal{N}\left(\mu,\Sigma\right)\quad\text{with }\mu=\begin{pmatrix}50\\ 20\\ 10\\ 100\end{pmatrix}\quad\Sigma=\begin{pmatrix}1&\text{Cov}(A,B)&\text{Cov}(A,C)&\text{Cov}(A,D)\\ \text{Cov}(A,B)&1&\text{Cov}(B,C)&\text{Cov}(B,D)\\ \text{Cov}(A,C)&\text{Cov}(B,C)&1&\text{Cov}(C,D)\\ \text{Cov}(A,D)&\text{Cov}(B,D)&\text{Cov}(C,D)&1\end{pmatrix}$ (30)
+The random variables $A,B,C,$ and $D$ are modeled as components of a four-dimensional Gaussian vector $(A,B,C,D)^{\top}\sim\mathcal{N}\left(\mu,\Sigma\right)$ with
+$\mu=\begin{pmatrix}50\\ 20\\ 10\\ 100\end{pmatrix}\quad\Sigma=\begin{pmatrix}1&\text{Cov}(A,B)&\text{Cov}(A,C)&\text{Cov}(A,D)\\ \text{Cov}(A,B)&1&\text{Cov}(B,C)&\text{Cov}(B,D)\\ \text{Cov}(A,C)&\text{Cov}(B,C)&1&\text{Cov}(C,D)\\ \text{Cov}(A,D)&\text{Cov}(B,D)&\text{Cov}(C,D)&1\end{pmatrix}$
 With these parameters, the ratio $R=\mathbb{E}[A]/\mathbb{E}[C]$ to be estimated equals 5.
@@
 Unless otherwise specified, the sample size $n$ is 100.
-Three particular configurations are considered in the following to illustrate possible behaviours of the estimators depending on the control variates coefficients.
+Three particular configurations are considered in the following to illustrate possible behaviors of the estimators depending on the control variates coefficients.
 The first corresponds to the best-case scenario for the optimal coefficients proposed in this work, i.e.
 the covariance structure that leads to the largest variance reduction.
-The second corresponds to the worst-case scenario for the coefficients of [Gordon et al., 1982], i.e.
+The second corresponds to the worst-case scenario for the coefficients of [gordon_efficient_1982], i.e.
 the covariance structure that leads to the largest variance increase.
-The last one corresponds to the best-case scenario for the coefficients of [Gordon et al., 1982].
+The last one corresponds to the best-case scenario for the coefficients of [gordon_efficient_1982].
 The covariance configurations are obtained by maximizing or minimizing the variance reduction using the formulas in A.2.
-The optimization is carried out using the differential evolution algorithm [Storn and Price, 1997], with the objective function set to infinity whenever the tested covariance matrix is not positive definite.
+The optimization is carried out using the differential evolution algorithm [storn_differential_1997], with the objective function set to infinity whenever the tested covariance matrix is not positive definite.
 4.1 Strong variance reduction
@@
 $\Sigma=\begin{pmatrix}1&&&\\ 0.27&1&&\\ -0.99&-0.29&1&\\ 0.95&-0.02&-0.95&1\end{pmatrix}$
-Figure 1: Boxplots of ratio of means estimations, comparing the MC/MC, CV/MC, and CV/CV estimators
-Strong variance reduction with all estimators
+Figure 1: Boxplots of ratio of means estimations, comparing the MC/MC, CV/MC, and CV/CV estimators - Strong variance reduction with all estimators
 As shown in Figure 1, the optimal coefficients $(\alpha_{o},\beta_{o})$ achieve greater variance reduction than alternative choices.
@@
 4.2 Variance increase with non-optimal coefficients
-The presented results are the worst-case scenario for the coefficients of [Gordon et al., 1982], happening with the covariance matrix
+The presented results are the worst-case scenario for the coefficients of [gordon_efficient_1982], happening with the covariance matrix
 $\Sigma=\begin{pmatrix}1&&&\\ -0.99&1&&\\ 0.99&-0.99&1&\\ -0.02&0.02&-0.01&1\end{pmatrix}$
-Figure 2: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators
-Variance increase with non-optimal coefficients
+Figure 2: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators - Variance increase with non-optimal coefficients
 While the variance reduction achieved with the optimal coefficients may not always exceed that of the other coefficients, their key advantage lies in their theoretical guarantee of variance reduction.
-The classical coefficients and the ones from [Gordon et al., 1982] can lead to a variance increase as illustrated in Figure 2.
+The classical coefficients and the ones from [gordon_efficient_1982] can lead to a variance increase as illustrated in Figure 2.
 This result highlights the importance of jointly optimizing the coefficients to minimize the variance of the whole estimator.
-4.3 Best variance reduction with the coefficients of [Gordon et al., 1982]
-The presented results are the best-case scenario for the coefficients of [Gordon et al., 1982], with the following covariance matrix
+4.3 Best variance reduction with the coefficients of [gordon_efficient_1982]
+Figure 3: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimator - Best variance reduction with the coefficients from [gordon_efficient_1982] with $n=10$
+Figure 4: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimator - Equivalent variance reduction performances with $n=100$
+The presented results are the best-case scenario for the coefficients of [gordon_efficient_1982], with the following covariance matrix
 $\Sigma=\begin{pmatrix}1&&&\\ -0.58&1&&\\ -0.99&0.57&1&\\ 0.99&-0.49&-0.99&1\end{pmatrix}$
-Figure 3: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators
-Best variance reduction with the coefficients from [Gordon et al., 1982] with $n=10$
 The optimal coefficients introduced in this work come with a theoretical guarantee of variance reduction and can outperform previously proposed coefficients.
@@
 Nevertheless, the differences in variance reduction are small and the optimal coefficients give the best results with the same covariance structure and more samples, as seen in Figure 4 with $n=100$.
-Figure 4: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators
-Equivalent variance reduction performances with $n=100$
-5 Application on a multifidelity aircraft design use case
-Figure 5: ONERA ALBATROS configuration with strut-braced wings (Figure 11 from [Carrier et al.,])
-Figure 6: Illustrations of the $1,252$ HF and LF samples of the dataset
+5 Applications
+5.1 Estimation of a proportion in a multifidelity aircraft design use case
 The CV/CV estimator is applied to estimate the strut mass fraction, defined as the ratio between the strut mass and the total mass.
-As illustrated in Figure 5, the strut is an external support between the fuselage and the wing that helps carry aerodynamic loads.
+The strut is an external support between the fuselage and the wing that helps carry aerodynamic loads.
 The strut mass fraction is a quantity of interest to be optimized as it ensures the added support does not negate the aerodynamic and structural advantages of a strut-braced wing.
-The studied dataset contains 1,252 samples from an aircraft design study, based on [Méheut et al., 2024] which describes the multidisciplinary and multi-fidelity design process.
-The dataset combines outputs from high fidelity (HF) [Lannoo and Priasso, 2025] and low fidelity (LF) [Carrier et al.,] models to efficiently explore the design space.
+The studied dataset contains 1,252 samples from an aircraft design study, based on [meheut_strut_braced_2024] which describes the multidisciplinary and multi-fidelity design process.
+The dataset combines outputs from high fidelity (HF) [lannoo_mass_2025] and low fidelity (LF) [carrier_multidisciplinary_2022] models to efficiently explore the design space.
 By applying the proposed CV/CV method, the strut mass fraction can be estimated more precisely with fewer HF model evaluations, which take around 2 minutes each, leveraging the LF model, which runs in only a few seconds and serves as a control variate as in MFMC.
@@
 $\begin{pmatrix}1&&&\\ 0.51&1&&\\ 0.77&0.4&1&\\ 0.83&0.78&0.74&1\end{pmatrix}$
+Figure 5: Illustrations of the $1,252$ coupled HF and LF samples of the aircraft design dataset
 (a) $n=200$ HF samples
@@
 In both scenarios, our approach achieves an RVR of about 20%, confirming its effectiveness in improving estimation accuracy and delivering the precision of a larger HF dataset at no extra computational cost.
+5.2 Estimation of a conditional value-at-risk in a multi-fidelity electromagnetic dataset
+A conditional Value-at-Risk (cVaR) [hong_monte_2014], also referred to as superquantile or expected shortfall [acerbi_coherence_2002] measures the average of values above a given quantile.
+While a quantile marks a threshold, the cVaR reflects the magnitude of extreme outcomes, providing a more robust assessment of risk widely applied in finance, reliability, and risk management [rockafellar_optimization_2000].
+The cVaR $q^{*}_{\alpha}$ of $X$ of order $\alpha\in[0,1]$ is defined as
+$q^{*}_{\alpha}(X)=\mathbb{E}[X|X\geq q_{\alpha}(X)]=\frac{\mathbb{E}[X\mathds{1}\{X\geq q_{\alpha}(X)\}]}{\mathbb{E}[\mathds{1}\{X\geq q_{\alpha}(X)\}]}$ (30)
+where $X\in L^{2}$ is an arbitrary random variable, with cumulative distribution function $F$, and $q_{\alpha}(X)=\text{inf}\{x\in\mathbb{R}|F_{X}(x)\geq\alpha\}$ is the quantile of $X$ of order $\alpha$.
+The studied multi-fidelity electromagnetic dataset is composed of $n=51$ HF samples $X^{HF}$ and $m=5,000$ additional LF samples $X^{LF}$.
+Aircraft certification for electromagnetic environments requires evaluating induced constraints, in terms of currents and voltages, at embedded equipment inputs.
+In the case of lightning indirect effects, the waveform rise time is a critical indicator, corresponding to the time in seconds between 10% and 90% of the induced current waveform.
+Following the notation previously used in this paper, the cVaR estimator can be written as the ratio $R=\mathbb{E}[A]/\mathbb{E}[C]$ where
+$A=X^{HF}\mathds{1}\{X^{HF}\geq q_{\alpha}(X^{HF})\}\qquad C=\mathds{1}\{X^{HF}\geq q_{\alpha}(X^{HF})\}$
+The control variates are defined analogously using the LF samples:
+$B=X^{LF}\mathds{1}\{X^{LF}\geq q_{\alpha}(X^{LF})\}\qquad D=\mathds{1}\{X^{LF}\geq q_{\alpha}(X^{LF})\}$
+Since their exact expectations are unknown, we apply the approximate control variates method, estimating their means using the full set of $n+m=5,051$ available samples.
+Due to the limited number of HF samples, bootstrapping is not appropriate, and the RVR is computed using the analytical formulas of the MC/MC and CV/CV estimators’ variance derived in Appendix A.3 in (34) and (38).
+(a) Illustration of the $51$ HF and LF coupled samples of the electromagnetic dataset
+(b) cVaR estimates for various quantile orders $\alpha$
+(c) RVR estimates for various quantile orders $\alpha$
+Figure 7: Results on the electromagnetic dataset
+As shown in Figure 7(a), the HF and LF samples are strongly correlated, suggesting potential variance reduction using the CV/CV estimator.
+Indeed, the estimated RVR depicted in Figure 7(c) ranges from 10% to 30%.
+At $\alpha=0.95$, non-optimal coefficients increase variance, and the variance reduction guarantee of the optimal coefficients is illustrated.
 6 Conclusion
@@
 In particular, we showed that previously proposed coefficients could lead to a variance increase.
-To illustrate the approach, we applied the estimators with approximate control variates to estimate the strut mass fraction on a multi-fidelity aircraft design dataset.
-Future work could aim to apply this method to reduce the variance of conditional mean estimators, such as the tail dependence [Malevergne and Sornette, 2002] or the conditional value at risk [Hong et al., 2014], as well as the self-normalized importance sampling estimator [Bugallo et al., 2017].
+To illustrate the approach, we applied the estimators with approximate control variates to estimate the strut mass fraction on a multi-fidelity aircraft design dataset and conditional values-at-risk on a multi-fidelity electromagnetic dataset.
+Future work could aim to apply this method to reduce the variance of conditional mean estimators, such as the tail dependence [malevergne_investigating_2002] or the self-normalized importance sampling estimator [bugallo_adaptive_2017].
 The method could also be generalized to estimators defined as functions of ratios.
 However, the current coefficients do not guarantee variance reduction in that setting; new coefficients would need to be derived for the full estimator.
+Acknowledgments
+We gratefully acknowledge Eric Nguyen Van (ONERA/DTIS) for providing the aircraft design dataset and sharing his insights into the associated design considerations, and Isabelle Junqua (ONERA/DEMR) for providing the electromagnetic dataset along with her guidance on the relevant physical considerations.
 Supplementary Materials
 All simulation codes used to produce the figures are available on GitHub at https://github.com/LouB-N/Control-variates-for-variance-reduced-ratio-of-means-estimators.git.
-Acknowledgements
-We gratefully acknowledge Eric Nguyen Van (ONERA/DTIS) for providing the dataset used in the application and for his valuable explanations of the aircraft design considerations behind it.
+Data Availability Statement
+The electromagnetic dataset and corresponding code to reproduce all figures are accessible on GitHub at https://github.com/LouB-N/Control-variates-for-variance-reduced-ratio-of-means-estimators.git.
+The aircraft design dataset is under consideration for publication; however, it can be provided to reviewers upon request.
+Disclosure statement
+The authors report there are no competing interests to declare.
+Declaration of generative AI use in the manuscript preparation process
+During the preparation of this work, the authors used Chat-GPT 5 in order to improve language and readability.
+After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
 Appendix A Deriving the optimal parameters and proof of variance reduction
@@
 Denote $X/Z$ the studied ratio, where $X$ and $Z$ are real random variables with non-zero means.
-Using first-order approximations and the delta method [Casella and Berger, 2024, example 5.5.27] [Sukhatme et al., 1970, section 4a.7] [Cochran, 1977, section 6.9], the variance of the Monte-Carlo estimator $\frac{\overline{X_{n}}}{\overline{Z_{n}}}$ of this ratio can be approximated by:
-$\boxed{\text{Var}\left(\frac{\overline{X_{n}}}{\overline{Z_{n}}}\right)=\frac{1}{n}\left(\frac{1}{\mathbb{E}[Z]^{2}}\text{Var}(X)+\frac{\mathbb{E}[X]^{2}}{\mathbb{E}[Z]^{4}}\text{Var}(Z)-2\frac{\mathbb{E}[X]}{\mathbb{E}[Z]^{3}}\text{Cov}\left(X,Z\right)\right)+\mathcal{O}(n^{-2})}$ (31)
+Using first-order approximations and the delta method [casella_properties_2024, example 5.5.27] [sukhatme1970sampling, section 4a.7] [cochran_sampling_1977, section 6.9], the variance of the Monte-Carlo estimator $\frac{\overline{X_{n}}}{\overline{Z_{n}}}$ of this ratio can be approximated by:
+$\boxed{\text{Var}\left(\frac{\overline{X_{n}}}{\overline{Z_{n}}}\right)=\frac{1}{n}\left(\frac{1}{\mathbb{E}[Z]^{2}}\text{Var}(X)+\frac{\mathbb{E}[X]^{2}}{\mathbb{E}[Z]^{4}}\text{Var}(Z)-2\frac{\mathbb{E}[X]}{\mathbb{E}[Z]^{3}}\text{Cov}\left(X,Z\right)\right)+\mathcal{O}(n^{-\frac{3}{2}})}$ (31)
 Following the same arguments, it can be shown that the approximation for the variance of the CV/MC and CV/CV estimators is of the same order.
@@
 Proposition 6 (Variance of the MC/MC estimator).
-Let $(A_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$ and $C$. The variance of the MC/MC estimator can be approximated as
+Let $(A_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A,C\in L^{2}$. The variance of the MC/MC estimator can be approximated as
 $\boxed{\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\approx\frac{1}{n\mathbb{E}[C]^{2}}\left(\text{Var}(A)+R^{2}\text{Var}(C)-2RCov(A,C)\right)}$ (34)
@@
 Let $(A_{i},B_{i},C_{i})_{i=1\dots n}$ be i.i.d.
-real samples from the joint distribution of the random variables $A$, $B$, and $C$, where $B$ is a control variate.
+real samples from the joint distribution of the random variables $A,B,C\in L^{2}$, where $B$ is a control variate.
 The variance of the CV/MC estimator can be approximated as
@@
 Let $(A_{i},B_{i},C_{i},D_{i})_{i=1\dots n}$ be i.i.d.
-real samples from the joint distribution of the random variables $A$, $B$, $C$, and $D$, where $B$ and $D$ are control variates.
+real samples from the joint distribution of the random variables $A,B,C,D\in L^{2}$, where $B$ and $D$ are control variates.
 The variance of the CV/CV estimator can be approximated as
@@
 Using the results from Lemma 3, the CV/CV estimator’s variance is derived as:
-$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$ $\approx\frac{1}{\mathbb{E}[\hat{C}_{CV,n}]^{2}}\text{Var}(\hat{A}_{CV})+\frac{\mathbb{E}[\hat{A}_{CV}]^{2}}{\mathbb{E}[\hat{C}_{CV,n}]^{4}}\text{Var}(\hat{C}_{CV,n})-2\frac{\mathbb{E}[\hat{A}_{CV}]}{\mathbb{E}[\hat{C}_{CV,n}]^{3}}\text{Cov}\left(\hat{A}_{CV},\hat{C}_{CV,n}\right)$
+$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$
+$\approx\frac{1}{\mathbb{E}[\hat{C}_{CV,n}]^{2}}\text{Var}(\hat{A}_{CV})+\frac{\mathbb{E}[\hat{A}_{CV}]^{2}}{\mathbb{E}[\hat{C}_{CV,n}]^{4}}\text{Var}(\hat{C}_{CV,n})-2\frac{\mathbb{E}[\hat{A}_{CV}]}{\mathbb{E}[\hat{C}_{CV,n}]^{3}}\text{Cov}\left(\hat{A}_{CV},\hat{C}_{CV,n}\right)$
 $\approx\frac{1}{\mathbb{E}[C]^{2}}\text{Var}(\hat{A}_{CV})+\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\hat{C}_{CV,n})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}\left(\hat{A}_{CV},\hat{C}_{CV,n}\right)$
@@
 $\nabla\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)=\begin{bmatrix}0\\ 0\end{bmatrix}\Leftrightarrow M\cdot\begin{bmatrix}\alpha_{o}\\ \beta_{o}\end{bmatrix}=\begin{bmatrix}c_{1}\\ c_{2}\end{bmatrix}$
-with $M=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})&-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\\ -2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})&2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})\end{bmatrix}$ and $\begin{bmatrix}c_{1}\\ c_{2}\end{bmatrix}=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})\\ 2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(\overline{C_{n}},\overline{D_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{D_{n}})\end{bmatrix}$.
+with $M=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})&-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\\ -2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})&2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})\end{bmatrix}$
+and $\begin{bmatrix}c_{1}\\ c_{2}\end{bmatrix}=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})\\ 2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(\overline{C_{n}},\overline{D_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{D_{n}})\end{bmatrix}$.
 The solution is given by: $\begin{bmatrix}\alpha_{o}\\ \beta_{o}\end{bmatrix}=M^{-1}\cdot\begin{bmatrix}c_{1}\\ c_{2}\end{bmatrix}$, where the inverse of $M$ is:
@@
 Note $\eta:=\frac{\text{Cov}(A,D)-R\text{Cov}(C,D)}{\text{Var}(D)}$. By replacing $\alpha$ and $\beta$ with their optimal values $\alpha_{o}$ and $\beta_{o}$, the variance expression is as follows:
-$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$ $\approx\frac{1}{n\mathbb{E}[C]^{2}}\left.\begin{aligned} \Big(&\text{Var}(A)+\left(\eta+\beta_{o}R\right)^{2}\text{Var}(D)-2\left(\eta+\beta_{o}R\right)\text{Cov}(A,D)+R^{2}\text{Var}(C)+\beta_{o}^{2}R^{2}\text{Var}(D)\\ &-2\beta_{o}R^{2}\text{Cov}(C,D)-2R\text{Cov}(A,C)+2\left(\eta+\beta_{o}R\right)R\text{Cov}(C,D)+2\beta_{o}R\text{Cov}(A,D)\\ &-2\left(\eta+\beta_{o}R\right)\beta_{o}R\text{Var}(D)\Big)\end{aligned}\right$.
+$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$
+$\approx\frac{1}{n\mathbb{E}[C]^{2}}\left.\begin{aligned} \Big(&\text{Var}(A)+\left(\eta+\beta_{o}R\right)^{2}\text{Var}(D)-2\left(\eta+\beta_{o}R\right)\text{Cov}(A,D)+R^{2}\text{Var}(C)+\beta_{o}^{2}R^{2}\text{Var}(D)\\ &-2\beta_{o}R^{2}\text{Cov}(C,D)-2R\text{Cov}(A,C)+2\left(\eta+\beta_{o}R\right)R\text{Cov}(C,D)+2\beta_{o}R\text{Cov}(A,D)\\ &-2\left(\eta+\beta_{o}R\right)\beta_{o}R\text{Var}(D)\Big)\end{aligned}\right$.
 $\approx\frac{1}{n\mathbb{E}[C]^{2}}\left.\begin{aligned} \Big(&\text{Var}(A)+\eta^{2}\text{Var}(D)-2\eta\text{Cov}(A,D)+R^{2}\text{Var}(C)-2R\text{Cov}(A,C)+2\eta R\text{Cov}(C,D)\\ &+2\eta\beta_{o}R\text{Var}(D)+\beta_{o}^{2}R^{2}\text{Var}(D)-2\beta_{o}R\text{Cov}(A,D)+\beta_{o}^{2}R^{2}\text{Var}(D)-2\beta_{o}R^{2}\text{Cov}(C,D)\\ &+2\beta_{o}R^{2}\text{Cov}(C,D)+2\beta_{o}R\text{Cov}(A,D)-2\eta\beta_{o}R\text{Var}(D)-2\beta_{o}^{2}R^{2}\text{Var}(D)\Big)\end{aligned}\right$.
@@
 There is a variance reduction if $\textbf{T}<0$.
-$\bullet$ With the coefficients from [Gordon et al., 1982] $\alpha_{g}=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ and $\beta_{g}=\frac{\text{Cov}(C,D)-\frac{1}{R}\text{Cov}(A,D)+\frac{\alpha}{R}\text{Cov}(B,D)}{\text{Var}(D)}$, the variance difference is:
+$\bullet$ With the coefficients from [gordon_efficient_1982] $\alpha_{g}=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ and $\beta_{g}=\frac{\text{Cov}(C,D)-\frac{1}{R}\text{Cov}(A,D)+\frac{\alpha}{R}\text{Cov}(B,D)}{\text{Var}(D)}$, the variance difference is:
 $\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$
@@
 $\text{Var}\left(\widehat{R}_{\frac{ACV}{ACV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)=\frac{m}{n+m}\left(\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)\right)$ (43)
-Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
-During the preparation of this work, the authors used ChatGPT in order to improve language and readability.
-After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.